### v1

# An Actionable Framework for Assessing Bias and Fairness in Large Language Model Use Cases

Abstract
Large language models (LLMs) can exhibit bias in a variety of ways. Such biases can create or exacerbate unfair outcomes for certain groups within a protected attribute, including, but not limited to sex, race, sexual orientation, or age. This paper aims to provide a technical guide for practitioners to assess bias and fairness risks in LLM use cases. The main contribution of this work is a decision framework that allows practitioners to determine which metrics to use for a specific LLM use case. To achieve this, this study categorizes LLM bias and fairness risks, maps those risks to a taxonomy of LLM use cases, and then formally defines various metrics to assess each type of risk. As part of this work, several new bias and fairness metrics are introduced, including innovative counterfactual metrics as well as metrics based on stereotype classifiers. Instead of focusing solely on the model itself, the sensitivity of both prompt-risk and model-risk are taken into account by defining evaluations at the level of an LLM use case, characterized by a model and a population of prompts. Furthermore, because all of the evaluation metrics are calculated solely using the LLM output, the proposed framework is highly practical and easily actionable for practitioners. [^1] Note: The examples provided in this paper are purely hypothetical and are not intended to reflect the specific work or practices of the author-affiliated company. They are used solely for illustrative purposes. Any resemblance to actual practices or projects is coincidental.
1 Introduction
The versatility of current Large Language Models (LLMs) in handling various tasks (Minaee et al., 2024; Liu et al., 2023; Ray, 2023) presents challenges when it comes to evaluating bias and fairness at the model level. Existing approaches primarily focus on assessing risk using benchmark data sets containing predefined prompts (Gehman et al., 2020; Dhamala et al., 2021; Nozza et al., 2021; Smith et al., 2022; Parrish et al., 2021; Li et al., 2020), masked tokens (Zhao et al., 2018; Rudinger et al., 2018; Nadeem et al., 2021; Levy et al., 2021), or unmasked sentences (Nangia et al., 2020; Barikeri et al., 2021; Jiao et al., 2023; Felkner et al., 2023), assuming that these adequately capture specific bias or fairness risks (Gallegos et al., 2023). However, these assessments are likely to overestimate the risk for use cases where the population of prompts is low risk (Wang et al., 2024). Moreover, to the best of the author’s knowledge, the current literature does not provide a framework for effectively aligning LLM use cases with suitable metrics for evaluating bias and fairness.
This work aims to address these limitations by developing an actionable LLM bias and fairness evaluation framework defined at the use case level. Drawing inspiration from the classification fairness framework proposed by Saleiro et al. (2018), the framework proposed in this work enables practitioners to map an LLM use case to an appropriate set of bias and fairness evaluation metrics by considering relevant characteristics of the use case and stakeholder values. This evaluation approach is unique in that it incorporates actual prompts from the practitioner’s use case, taking into account the prompt-specific risks that have been demonstrated to significantly increase the likelihood of biased and unfair outcomes (Wang et al., 2024). By constraining the scope to focused use cases, where prompts are derived from a known population and the task is well-defined, this framework is specifically designed to customize the risk assessment for a specific application.
To introduce the framework, this study first provides formal definitions of bias and fairness desiderata for LLMs from the literature and segment these definitions by risk category. Subsequently, those risks are mapped to a taxonomy of use cases focused on large-scale applications, where human-in-the-loop may be infeasible. Lastly, for each risk category, various bias and fairness evaluation metrics are detailed, with discussions provided on their input requirements, calculation methods, the risks they assess, and circumstances under which they should be applied. As part of this work, a variety of novel bias and fairness metrics are introduced. This includes innovative counterfactual adaptations of recall-oriented understudy for gisting evaluation (ROUGE) (Lin, 2004), bilingual evaluation understudy (BLEU) (Papineni et al., 2002), and cosine similarity (Singhal and Google, 2001), as well a set of stereotype classifier-based metrics that are adapted from analogous toxicity classifier-based metrics.
For practical reasons, this study limits the selection of LLM bias and fairness metrics to those requiring only LLM generated output as inputs. This includes 1) generated text metrics, which take a generated set of tokens as input (Gallegos et al., 2023), 2) recommendation fairness metrics, calculated on a set of LLM-provided recommendations (Zhang et al., 2023), and 3) classification fairness metrics, which are already well-established in the machine learning fairness literature (Bellamy et al., 2018; Saleiro et al., 2018; Weerts et al., 2023; Hardt et al., 2016; Feldman et al., 2014; Mehrabi et al., 2019). Due to practical limitations of the input requirements, the framework omits both embedding-based metrics, which are computed using an LLM’s hidden vector representations of words or sentences (Islam et al., 2016; May et al., 2019; Guo and Caliskan, 2020), and probability-based metrics, which leverage predicted token probabilities from an LLM (Webster et al., 2020; Kurita et al., 2019; Ahn and Oh, 2021; Kaneko and Bollegala, 2021; Salazar et al., 2019; Barikeri et al., 2021; Nangia et al., 2020; Nadeem et al., 2021). To further support the rationale behind the selection of evaluation metrics, it is important to note that metrics focused on the downstream task, consistent with the metrics incorporated in this framework, have been shown to be more reliable than metrics derived from embeddings or token probabilities (Goldfarb-Tarrant et al., 2020; Delobelle et al., 2021).
The remainder of this paper is organized as follows. Section 2 formally defines various notions of bias and fairness in LLMs. Section 3 discusses various techniques for conducting bias and fairness assessments in LLMs. Section 4 offers a framework for choosing among LLM bias and fairness metrics based on use case characteristics and stakeholder values. Finally, Section 5 offers concluding remarks.
2 Bias and Fairness Risks for LLM Use Cases
This section formally defines several pre-requisite terms and concepts upon which subsequent sections rely, several of which are adapted from those provided by Gallegos et al. (2023). [^2] The notation and terminology used in this paper differs from those used by Gallegos et al. (2023). The approach outlined below adopts modified versions of these definitions to better fit the tone and purpose. For further reading on biases in LLMs, the reader may refer to (Gallegos et al., 2023; Blodgett et al., 2020; Kumar et al., 2023; Li et al., 2024; Chu et al., 2024; Ferrara, 2023; Ranaldi et al., 2023; Kotek et al., 2023; Wu and Aji, 2023; Li et al., 2023; Ray, 2023; Nozza et al., 2022; Zhuo et al., 2023). The discussions in this paper explore concepts of bias and fairness in relation to an arbitrary ‘protected attribute’, encompassing examples such as sex, race, age, and sexual orientation, among others. These concepts are established formally by using mathematical notation whenever possible and strives to prioritize consistency in notation throughout, even if it may differ from the notation used in the papers cited.
2.1 Preliminary Definitions
This section provides preliminary definitions to be used throughout the subsequent sections.
Large Language Model (LLM).
An LLM $\mathcal{M}:\mathcal{X}\xrightarrow{}\mathcal{Y}$ is a pre-trained, transformer-based model that maps a text sequence $X\in\mathcal{X}$ to an output $\hat{Y}\in\mathcal{Y}$, where $\mathcal{X}$ denotes the set of all possible text inputs (i.e. prompts) and the form of $\hat{Y}$ is specific to the LLM and the use case (Gallegos et al., 2023). [^3] Following Gallegos et al. (2023), $\mathcal{M}$ is assumed to have an autoregressive, autoencoding, or encoder-decoder architecture and has undergone training on an extensive corpus containing hundreds of millions to trillions of tokens.  [^4] Hereafter, all LLM inputs $X$ are referred to as ‘prompts’. Let $\mathcal{X}$ denote the set of all possible prompts. Let $\theta$ parameterize $\mathcal{M}$, such that $\hat{Y}=\mathcal{M}(X;\theta)$.
Population of Prompts.
A population of prompts, denoted $\mathcal{P}_{X}$, is a collection of LLM inputs. To characterize well-defined use cases, subsequent sections refer to a ‘known population of prompts’, indicating that practitioners possess information about the prompt domain and are able draw representative samples from $\mathcal{P}_{X}$. For instance, a population of prompts might consist of clinical notes, where each individual prompt includes a collection of notes, accompanied by specific instructions for the LLM to generate a summary (Chuang et al., 2024).
Large Language Model Use Case.
An LLM use case is characterized by an LLM $\mathcal{M}(X;\theta)$ and a population of prompts $\mathcal{P}_{X}$. In the interest of concise notation, LLM use cases will be hereafter denoted as $(\mathcal{M},\mathcal{P}_{X}).$ An LLM use case is evaluated on a finite set of responses generated by $\mathcal{M}(X;\theta)$ from a sample of $N$ prompts $X_{1},...,X_{N}$, drawn from the population $\mathcal{P}_{X}$.
Protected Attribute Groups.
A protected attribute group $G\in\mathcal{G}$ represents a subset of people characterized by a shared identity trait, where $\mathcal{G}$ is a partition (Gallegos et al., 2023).
Protected Attribute Group Lexicon.
A protected attribute group lexicon $A\in\mathcal{A}$ is a collection of words that correspond to protected attribute group $G\in\mathcal{G}$ [^5] Hereafter, in a slight abuse of notation, $\mathcal{G}$ (and analogously $\mathcal{A}$) can denote either a partition of or the union of each $G\in\mathcal{G}$ (and analogously each $A\in\mathcal{A}$). [^6] One example, provided by Bommasani et al. (2023), of protected attribute words for the protected attribute group ‘males’ is { ‘he’, ‘son’, ‘his’, ‘him’, ‘father’, ‘man’, ‘boy’, ‘himself’, ‘male’, ‘brother’, ‘sons’, ‘fathers’, ‘men’, ‘boys’, ‘males’, ‘brothers’, ‘uncle’, ‘uncles’, ‘nephew’, ‘nephews’ }.  [^7] The examples that consider male vs. female are used in accordance with the convention established by previous studies such as Bordia and Bowman (2019) and Bommasani et al. (2023). It is important to note that this usage is solely for the purpose of consistency and does not intend to imply that there are only two genders. The acknowledgement of a broader spectrum of gender identities is recognized and respected.
Counterfactual Input Pair.
[…]
Fairness Through Unawareness (FTU).
Given a protected attribute lexicon $\mathcal{A}$, an LLM use case $(\mathcal{M},\mathcal{P}_{X})$ satisfies FTU if for each $X\in\mathcal{P}_{X},X\cap\mathcal{A}=\emptyset$. In simpler terms, FTU implies none of the prompts for an LLM use case include any mention of a protected attribute word (Gallegos et al., 2023).
2.2 LLM Bias and Fairness Risks
This section presents formal definitions of various notions of bias and fairness applicable to LLMs. A taxonomy for these definitions is proposed, organized by risk category. Specifically, the risk categories included in this taxonomy are toxicity, stereotyping, counterfactual fairness, and allocational harms.
2.2.1 Toxicity
Drawing from the definitions of toxicity and derogatory language outlined in the survey conducted by Gallegos et al. (2023), let the definition of toxic text encompass any offensive language that 1) launches attacks, issues threats, or incites hate or violence against a social group, or 2) includes the usage of pejorative slurs, insults, or any other forms of expression that specifically target and belittle a social group. To formalize this, a corresponding fairness desideratum is introduced below.
Non-Toxicity.
Let $\mathcal{T}$ denote the set of all toxic phrases. An LLM use case $(\mathcal{M},\mathcal{P}_{X})$ exhibits non-toxicity if $\mathcal{M}(X;\theta)\cap\mathcal{T}=\emptyset$ for each $X\in\mathcal{P}_{X}$.
2.2.2 Stereotyping
Stereotyping is an important type of social bias that should be considered in the context of LLMs (Bommasani et al., 2023; Bordia and Bowman, 2019; Zekun et al., 2023). This study follows the work by Gallegos et al. (2023), in which stereotypes are defined as negative generalizations about a protected attribute group, often reflected by differences in frequency with which various groups are linked to stereotyped terms (Bommasani et al., 2023). To formalize this notion, two fairness desiderata are considered, which are also proposed by Gallegos et al. (2023): equal group associations (EGA) and equal neutral associations (ENA).
Equal Group Associations (Gallegos et al., 2023).
For two protected attribute groups $G^{\prime},G^{\prime\prime},$ and a set of neutral words $W$, an LLM use case $(\mathcal{M},\mathcal{P}_{X})$ satisfies equal group associations if, for each $w\in W,P(w\in\hat{Y}|\hat{Y}\cap A^{\prime}\neq\emptyset)=P(w\in\hat{Y}|\hat{Y}\cap A^{\prime\prime}\neq\emptyset)$. Put simply, equal group associations requires that each neutral word in $W$ is equally likely to be contained in the output of $\mathcal{M}$ regardless of which protected attribute group is mentioned.
Equal Neutral Associations (Gallegos et al., 2023).
For two protected attribute groups $G^{\prime},G^{\prime\prime},$ with respective associated lexicons $A^{\prime},A^{\prime\prime},$ and a set of neutral words $W$, an LLM use case $(\mathcal{M},\mathcal{P}_{X})$ satisfies Equal Neutral Associations if, $P(A^{\prime}\cap\hat{Y}\neq\emptyset|\hat{Y}\cap W\neq\emptyset)=P(A^{\prime\prime}\cap\hat{Y}\neq\emptyset|\hat{Y}\cap W\neq\emptyset).$ In other words, Equal Neutral Associations requires that co-occurrence of protected attribute words with a set of neutral words $W$ is equally probable for both protected attribute groups. [^9] In practice, the set $W$ referenced in the definitions of equal group associations and equal neutral associations typically contains frequently stereotyped terms such as professions or adjectives (Bordia and Bowman, 2019; Bommasani et al., 2023).
2.2.3 Counterfactual Fairness
In many contexts, it is undesirable for an LLM to generate substantially different output as a result of different protected attribute words contained in the input prompts, all else equal (Huang et al., 2020; Nozza et al., 2021; Wang et al., 2024). Following previous work (Huang et al., 2020; Garg et al., 2019), this concept is hereafter referred to as (lack of) counterfactual fairness. Depending on context and stakeholder values, the practitioner may wish to assess an LLM for differences in overall content or sentiment resulting from inclusion of different protected attribute words in a prompt. Below, this section provides a formal definition of the corresponding fairness desiderata known as counterfactual invariance, adapted from Gallegos et al. (2023).
Counterfactual Invariance.
For two protected attribute groups $G^{\prime},G^{\prime\prime},$ an LLM use case $(\mathcal{M},\mathcal{P}_{X})$ satisfies counterfactual invariance if, for a specified invariance metric $\upsilon(\cdot,\cdot)$, expected value of the invariance metric is less than some tolerance level $\epsilon$:
$\mathbb{E}[\upsilon(\mathcal{M}(X^{\prime};\theta),\mathcal{M}(X^{\prime\prime};\theta))]\leq\epsilon,$
where $(X^{\prime},X^{\prime\prime})$ is a counterfactual input pair corresponding to $G^{\prime},G^{\prime\prime}$ (Gallegos et al., 2023). [^10] This modified definition of counterfactual invariance relaxes strict equality as in Gallegos et al. (2023).
2.2.4 Allocational Harms
Allocational harms, which Gallegos et al. (2023) define as an unequal distribution of resources or opportunities among different protected attribute groups, have been widely studied in the machine learning fairness literature (Saleiro et al., 2018; Bellamy et al., 2018; Weerts et al., 2023; Kamishima et al., 2012; Zhang et al., 2018; Hardt et al., 2016; Feldman et al., 2014; Pleiss et al., 2017; Kamiran et al., 2012; Agarwal et al., 2018; Kamiran and Calders, 2011; Chouldechova, 2016). The corresponding fairness desideratum, is known as group fairness, defined formally below.
Group Fairness.
Given two protected attribute groups $G^{\prime},G^{\prime\prime},$ and a tolerance level $\epsilon$, an LLM use case $(\mathcal{M},\mathcal{P}_{X})$ satisfies group fairness if
$|B(\mathcal{M}(X;\theta)|G^{\prime})-B(\mathcal{M}(X;\theta)|G^{\prime\prime})|\leq\epsilon,$
where $B$ is a statistical performance metric (e.g. false negative rate) applied to $\mathcal{M}$, conditioned on membership in a protected attribute group (Gallegos et al., 2023). Here, conditioning on $G$ implies calculating $B$ on the subset of input prompts that either contain a direct mention of group $G$ or, in the case of person-level prompt granularity, correspond to individuals belonging to group $G$. Note that the choice of $B$ will depend on context and stakeholder values.
2.3 Mapping Bias and Fairness Risks to LLM Use Cases
Section 2.1 establishes the characterization of a well-defined LLM use case based on a model and a known population of prompts. [^11] Note that use case segmentation proposed in this work is by task, which can be controlled through various means such as fine-tuning of the LLM, providing examples with few-shot prompting, or incorporating instructions in system or user prompts. To set up the decision framework, this section segments use cases by task, according to the following three categories: 1) text generation and summarization, 2) classification, and 3) recommendation. Descriptions and examples are provided in Table 1.
The bias and fairness evaluation framework proposed in this work is intended for large-scale applications, in which the volume of generated responses makes exhaustive human review impractical. It is important to note that, for scenarios in which the practitioner manually evaluates each generated output, the evaluations proposed here may be unnecessary if concerns related to bias and fairness can be effectively addressed by the individual who is reviewing the outputs.
Table 1: Taxonomy of LLM Use Cases and Associated Bias/Fairness Risks
Use Case Category Description Examples Bias/Fairness Risk
Text Generation and Summarization LLM generates text outputs that are not constrained to a predefined set of classes or list elements Create personalized outreach messages to individuals; Summarize clinical notes Toxic text, stereotypes*, counterfactual fairness*
Classification LLM classifies a text input among a pre-defined set of classes Classify intent of customer support inquiries to assign assistance; Classify customer feedback as positive or negative to assign follow-ups Allocational harms**
Recommendation LLM generates lists of recommendations Generate lists of recommended products; Generate lists of recommended news articles Counterfactual fairness*
*Risk is applicable if FTU is not satisfied. Counterfactual fairness may not be relevant in certain contexts.
**Risk is applicable if text inputs correspond to a protected attribute.
Text Generation and Summarization.
First, consider use cases where an LLM generates text outputs that are not constrained to a predefined set of classes (e.g., positive vs. negative) or list elements (e.g., products to recommend). For the sake of brevity, this group of use cases will be hereafter referred to as “text generation and summarization," acknowledging that this category can encompass additional use cases including, but not limited to, machine translation, retrieval augmented generation (RAG), and question-answering. An example of a text generation use case could be utilizing an LLM to compose personalized messages for customer outreach. Similarly, a summarization use case may involve employing an LLM to extract pertinent information or provide summaries from clinical notes. These use cases carry the potential risk of generating toxic text in their outputs. Moreover, if these use cases fail to meet the criteria of FTU, meaning that the prompts include references to protected attributes, they also pose the risk of perpetuating stereotypes or exhibiting counterfactual unfairness.
Classification.
LLMs have been widely used for text classification (Sun et al., 2023; Widmann and Wich, 2022; Bonikowski et al., 2022; Howard and Ruder, 2018; Sun et al., 2019; Chai et al., 2020; Chen et al., 2020; Lin et al., 2021). In the context of bias and fairness, it is important to distinguish whether the text inputs can be mapped to a protected attribute, either by containing direct mentions of a protected attribute group, or in the case of person-level prompt granularity, corresponding to individuals belonging to certain protected attribute groups. For instance, one example of a person-level classification use case could involve utilizing an LLM to classify customer feedback as positive or negative in order to assign appropriate follow-ups. Similar to traditional person-level classification challenges in machine learning, these use cases present the risk of allocational harms. On the other hand, classification use cases that do not involve person-level data and satisfy FTU are not subject to these bias and fairness risks.
Recommendation.
Recommendation is another potential application of LLMs (Bao et al., 2023; Gao et al., 2023), such as using an LLM to recommend products to customers. Zhang et al. (2023) show that LLMs used as recommendation engines can discriminate when exposed to protected attribute information. Given this concern, it follows that LLM recommendation use cases pose the risk of counterfactual unfairness if they do not satisfy FTU.
3 Bias and Fairness Evaluation Metrics
This section introduces a range of evaluation metrics segmented by the use case task and its applicable bias and fairness risks. The proposed framework encompasses three distinct use case categories: 1) text generation and summarization, 2) classification, and 3) recommendation. For each category, bias and fairness evaluation metrics are included that address the applicable risks.
Accounting for prompt-specific risk is important to accurately reflect the risk of a specific use case, as Wang et al. (2024) find in their evaluations that toxicity probability is 26 to 101 times higher when dealing with toxic prompts compared to non-toxic prompts. Moreover, Goldfarb-Tarrant et al. (2020); Delobelle et al. (2021) have demonstrated that evaluation metrics which take into account the LLM’s task provide a more accurate reflection of the associated risk compared to metrics based on embeddings or token probabilities. Accordingly, the objective of this work is to assess the risks associated with a particular use case, taking into consideration not only the LLM used but also the task at hand and the prompt population. Following the characterization of a use case based on a model and a known population of prompts, as established in Section 2.1, each metric definition presented below is contextualized within an evaluation sample of size $N$ drawn from a known population of prompts $\mathcal{P}_{X}$.
3.1 Bias Metrics for Text Generation and Summarization Use Cases
Gallegos et al. (2023) propose a detailed taxonomy of evaluation metrics for bias evaluations in LLMs. They partition these metrics into three main categories: embedding-based metrics, probability-based metrics, and generated-text metrics. While embedding- and probability-based metrics require access to an LLM’s upstream architecture, generated-text metrics instead treat an LLM like a black box and can be easily calculated from LLM output alone. Hence, given their ease of application, only generated-text metrics are included in the framework proposed here.
While Gallegos et al. (2023) further segment generated-text metrics into three subcategories (namely distribution metrics, classifier metrics, and lexicon metrics), this work segments these metrics according to the risk taxonomy outlined in 2.2. In particular, generated-text bias and fairness metrics are segmented into the following categories: toxicity metrics, stereotype metrics, and counterfactual fairness metrics. Toxicity metrics leverage a pre-trained toxicity classifier, such as Perspective API [^12] https://perspectiveapi.com, to assign a toxicity score to an LLM’s output (Chowdhery et al., 2022; Lees et al., 2022; Wang et al., 2024; Bommasani et al., 2023; Gehman et al., 2020). Stereotype metrics assess the relative co-occurrence of stereotype words with protected attribute words (Bordia and Bowman, 2019; Bommasani et al., 2023) or leverage a pre-trained stereotype classifier to assign a stereotype score to an LLM’s output (Zekun et al., 2023). Lastly, counterfactual fairness metrics assess fairness through a causal lens by exploiting variation in mentions of protected attributes in otherwise-identical prompts (Huang et al., 2020). Select metrics are detailed in each category below. [^13] Note that the discussion of generated text bias metrics in this paper is not exhaustive. Rather, the intention is to select an appropriate set of evaluation metrics to address each risk to be assessed. Notable metrics omitted from this framework include HONEST (Nozza et al., 2021), score parity (Sicilia and Alikhani, 2023), Full Gen Bias (Smith et al., 2022), regard score (Sheng et al., 2019), demographic representations (Bommasani et al., 2023), psycholinguistic norms (Dhamala et al., 2021), and gender polarity (Dhamala et al., 2021).
3.1.1 Toxicity Metrics
Toxicity metrics leverage a pre-trained toxicity classifier to assess general toxicity that is not specific to a particular protected attribute. Given a toxicity classifier $T:\mathcal{Y}\xrightarrow[]{}[0,1]$, this section defines toxicity metrics in the context of an LLM use case $(\mathcal{M},\mathcal{P}_{X})$. Specifically, these metrics are evaluated on a set of $m$ responses per prompt, $\{\hat{Y}_{ij}\}_{j=1}^{m}$, generated by $\mathcal{M}$, from a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$. In practice, variation in responses for the same prompt can be achieved by choosing a non-zero temperature for the LLM or by utilizing various random seeds (Wang et al., 2024). Below, formal definitions of three toxicity metrics are provided: Expected Maximum Toxicity, Toxicity Probability, and Toxic Fraction. All three metrics range in values from 0 to 1, with smaller values indicating a greater degree of fairness.
Expected Maximum Toxicity (EMT) (Gehman et al., 2020).
EMT estimates the maximum predicted toxicity probability among the top $m$ generations:
$EMT=\frac{1}{N}\sum_{i=1}^{N}\max_{1\leq j\leq m}T(\hat{Y}_{ij}).$
Note that the standard choice of $m$ for this metric is $m=25$ (Gehman et al., 2020; Wang et al., 2024).
Toxicity Probability (TP) (Gehman et al., 2020).
TP is calculated as the empirical probability of having at least one toxic prediction ($T(\hat{Y})\geq 0.5$) among the top $m$ generations:
$TP=\frac{1}{N}\sum_{i=1}^{N}I(\max_{1\leq j\leq m}T(\hat{Y}_{ij})\geq 0.5).$
As with EMT, the standard choice of $m$ for this metric is $m=25$ (Gehman et al., 2020; Wang et al., 2024).
Toxicity Fraction (TF) (Bommasani et al., 2023).
TF is defined as the fraction of generations that are classified as toxic:
$TF=\frac{1}{N}\frac{1}{m}\sum_{i=1}^{N}\sum_{j=1}^{m}I(T(\hat{Y}_{ij})\geq 0.5),$
This metric effectively estimates the likelihood that responses generated by $\mathcal{M}$ on prompts from $\mathcal{P}_{X}$ contain toxic text (Bommasani et al., 2023). Note that while the standard choice of $m$ for this metric is $m=1$ (Bommasani et al., 2023), a larger value of $m$ may be preferred in practice if sampling a large $N$ is infeasible.
3.1.2 Stereotype Metrics
Stereotype metrics aim to identify harmful stereotypes specific to protected attributes that might be present in an LLM’s output. Because these metrics rely on mentions of protected attribute groups, these metrics may be unnecessary if FTU is satisfied for an LLM use case. [^14] Note that stereotype risk, while low, may still exist even if FTU is satisfied for an LLM use case. Among stereotype metrics, this work distinguishes between metrics based on co-occurrence of protected attribute words and stereotypical words, and metrics that leverage a stereotype classifier.
3.1.2.1 Co-occurrence-Based Metrics
This section outlines a set of metrics that assess stereotype risk based on relative co-occurrence of protected attribute words with neutral words of interest. In particular, formal definitions of two co-occurrence-based stereotype metrics are provided: Co-Occurrence Bias Score and Stereotypical Associations.
Co-Occurence Bias Score (COBS) (Bordia and Bowman, 2019).
Given two protected attribute groups $G^{\prime},G^{\prime\prime}$ with associated sets of protected attribute words $A^{\prime},A^{\prime\prime}$, a set of stereotypical words $W$, and an LLM use case $(\mathcal{M},\mathcal{P}_{X})$, the full calculation of COBS is as follows:
$P(w|A)=\frac{\sum_{i=1}^{N}cooccur(w,A|\hat{Y}_{i})/\sum_{i=1}^{N}\sum_{\tilde{w}\in\hat{Y}_{i}}cooccur(\tilde{w},A|\tilde{Y}_{i})}{\sum_{i=1}^{N}\sum_{a\in A}C(a,\hat{Y}_{i})/\sum_{i=1}^{N}\sum_{\tilde{w}\in\tilde{Y}_{i}}C(\tilde{w},\hat{Y}_{i})}$
$COBS=\frac{1}{|W|}\sum_{w\in W}\log\frac{P(w|A^{\prime})}{P(w|A^{\prime\prime})}.$
Above, $C(x,\hat{Y}_{i})$ denotes the count of $x$ in $\hat{Y}_{i}$ and $\tilde{Y}_{i}$ represents the LLM output $\hat{Y}_{i}$, with words from $A^{\prime}\cup A^{\prime\prime}$ and stop words excluded. The co-occurrence function $cooccur(w,A|\hat{Y})$ computes a weighted count of words from $A$ that are found within a context window centered around $w$, each time $w$ appears in $\hat{Y}$. [^15] To specify the co-occurrence function $cooccur(w,A|\hat{Y})$, Bordia and Bowman (2019) define a fixed or infinite context window centered at a target word $w$. In the case of fixed context (Bordia and Bowman (2019) recommend including 10 words before and 10 words after), each word in the context window receives equal weight. In the case of an infinite context window, each word receives weight $\beta^{dist},$ where $dist$ denotes the number of tokens between the two tokens of interest (Bordia and Bowman (2019) use $\beta=0.95$). In both cases, only words from the context window included in $A$ are included in the weighted count. In words, COBS computes the relative likelihood that an LLM $\mathcal{M}$ generates output having co-occurrence of $w\in W$ with $A^{\prime}$ versus $A^{\prime\prime}$. This metric has a range of possible values of [,][$\infty$,$\infty$][ - ∞ , ∞ ], with values closer to 0 signifying a greater degree of fairness.
Stereotypical Associations (SA) (Bommasani et al., 2023).
Consider a set of protected attribute groups $\mathcal{G}$, an associated set of protected attribute lexicons $\mathcal{A}$, and an associated set of stereotypical words $W$. Additionally, let $C(x,\hat{Y})$ denote the number of times that the word $x$ appears in the output $\hat{Y}$, $I(\cdot)$ denote the indicator function, $P^{\text{ref}}$ denote a reference distribution, and $TVD$ denote total variation difference. [^16] The reference distribution recommended by Bommasani et al. (2023) is the uniform distribution. Total variation distance measures the distance between probability distributions. For a given LLM $\mathcal{M}(X;\theta)$ and a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$, the full computation of SA is as follows:
$\gamma{(w|A^{\prime})}=\sum_{a\in A^{\prime}}\sum_{i=1}^{N}C(a,\hat{Y}_{i})I(C(w,\hat{Y}_{i})>0)$
$\pi(w|A^{\prime})=\frac{\gamma(w|A^{\prime})}{\sum_{A\in\mathcal{A}}\gamma(w|A)}$
$P^{(w)}=\{\pi(w|A^{\prime}):A^{\prime}\in\mathcal{A}\}$
$SA=\frac{1}{|W|}\sum_{w\in W}TVD(P^{(w)},P^{\text{ref}}).$
In words, SA measures the relative co-occurrence of a set of stereotypically associated words across protected attribute groups. SA ranges in value from 0 to 1, where smaller values indicate greater fairness.
3.1.2.2 Metrics Leveraging a Stereotype Classifier
It has been shown that stereotype classifiers can be an effective tool for assessing stereotype risk in LLM use cases (Zekun et al., 2023). [^17] Zekun et al. (2023) refer to https://huggingface.co/Narrativa/distilroberta-finetuned-stereotype-detection. This section introduces three new metrics by extending the toxicity metrics outlined in Section 3.1.1, leveraging a pre-trained stereotype classifier $St:\mathcal{Y}\xrightarrow[]{}[0,1]$ rather than a toxicity classifier. Namely, these metrics include: Expected Maximum Stereotype, Stereotype Probability, and Stereotype Fraction. All three metrics range in values from 0 to 1, with smaller values indicating a greater degree of fairness.
Expected Maximum Stereotype (EMS).
EMS, analogous to EMT, estimates the maximum predicted stereotype probability among the top $m$ generations:
$EMS=\frac{1}{N}\sum_{i=1}^{N}\max_{1\leq j\leq m}St(\hat{Y}_{ij}).$
Following the convention of EMT, practitioners may wish to use $m=25$ for this metric.
Stereotype Probability (SP).
Analogous to TP, SP measures as the empirical probability of having at least one stereotype prediction ($St(\hat{Y})\geq 0.5$), among the top $m$ generations:
$SP=\frac{1}{N}\sum_{i=1}^{N}I(\max_{1\leq j\leq m}St(\hat{Y}_{ij})\geq 0.5),$
To be consistent with the convention of TP, practitioners may wish to use $m=25$ for this metric.
Stereotype Fraction (SF).
SF, presented as an extension of TF, measures as the fraction of generations that are predicted to contain a stereotype:
$SF=\frac{1}{N}\frac{1}{m}\sum_{i=1}^{N}\sum_{j=1}^{m}I(St(\hat{Y}_{ij})\geq 0.5),$
effectively estimating the likelihood that responses generated by $\mathcal{M}$ on prompts from $\mathcal{P}_{X}$ contain stereotypes. Note that while the standard choice of $m$ for the analogous toxicity metric, EMT, is $m=1$ (Bommasani et al., 2023), a larger value of $m$ may be preferred in practice if sampling a large $N$ is infeasible.
3.1.3 Counterfactual Fairness Metrics
Counterfactual metrics aim to assess differences in LLM output when different protected attributes are mentioned in input prompts, all else equal. Given two protected attribute groups $G^{\prime},G^{\prime\prime}$, these metrics are defined in the context of an LLM use case $(\mathcal{M},\mathcal{P}_{X})$. In particular, these metrics are evaluated on a sample of counterfactual response pairs $(\hat{Y}_{1}^{\prime},\hat{Y}_{1}^{\prime\prime}),...,(\hat{Y}_{N}^{\prime},\hat{Y}_{N}^{\prime\prime})$ generated by $\mathcal{M}$, from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ drawn from $\mathcal{P}_{X|\mathcal{A}}$. [^18] In practice, counterfactual substitution can be achieved by leveraging a mapping of one protected attribute group lexicon to another. For instance, a female to male lexicon mapping could include substitutions such as ’she’: ’he’, ’hers’: ’his’, ’her’: ’him’, ’herself’: ’himself’, ’female’: ’male’, ’females’: ’males’, ’woman’: ’man’, ’women’: ’men’, ’girl’: ’boy’, ’girls’: ’boys’, ’daughter’: ’son’, ’daughters’: ’sons’, ’mother’: ’father’, ’mothers’: ’fathers’, ’sister’: ’brother’, ’sisters’: ’brothers’, ’aunt’: ’uncle’, ’aunts’: ’uncles’, ’niece’: ’nephew’, ’nieces’: ’nephews’, ’lady’: ’gentleman’, ’ladies’: ’gentlemen’, ’grandmother’: ’grandfather’, ’grandmothers’: ’grandfathers’. It is important to note that the mapping does not need to be exhaustive, as long as it has sufficient coverage to generate a large sample of counterfactual input pairs. Note that, in scenarios where a large $N$ is infeasible, practitioners may opt to generate multiple response pairs per counterfactual input pair.
These metrics, which are categorized into counterfactual similarity metrics and counterfactual sentiment metrics, respectively quantify the differences in text similarity and sentiment by leveraging the variations in LLM output observed across counterfactual input pairs. Due to their reliance on mentions of protected attributes in input prompts, if FTU is satisfied for an LLM use case, these metrics need not be used.
3.1.3.1 Counterfactual Similarity
Counterfactual similarity metrics measure the similarity in outputs generated from counterfactual input pairs according to a specified invariance metric $\upsilon$, i.e. $\upsilon(\mathcal{M}(X^{\prime};\theta),\mathcal{M}(X^{\prime\prime};\theta)).$ These metrics effectively assess whether the LLM use case satisfies the counterfactual invariance property defined in 2.2. One such example of $\upsilon$ is exact match (Rajpurkar et al., 2016), but (Gallegos et al., 2023) argue that this metric is too strict. This work introduces three, less stringent, counterfactual similarity metrics: Counterfactual ROUGE-L, Counterfactual BLEU, and Counterfactual Cosine Similarity, which are extensions of state-of-the-art text similarity metrics (Minaee et al., 2024; Lin, 2004; Papineni et al., 2002; Singhal and Google, 2001; Gomaa and Fahmy, 2013). While the first two assess similarity using token-sequence overlap, the the third assesses similarity using sentence embeddings. All three metrics range in values from 0 to 1, with larger values indicating a greater degree of fairness.
Counterfactual ROUGE-L (CROUGE-L).
This work introduces CROUGE-L, defined as the average ROUGE-L score (Lin, 2004) over counterfactually generated output pairs. The full calculation of CROUGE-L is as follows:
$r_{i}^{\prime}=\frac{LCS(\hat{Y}_{i}^{\prime},\hat{Y}_{i}^{\prime\prime})}{len(\hat{Y}_{i}^{\prime})}$
$r_{i}^{\prime\prime}=\frac{LCS(\hat{Y}_{i}^{\prime\prime},\hat{Y}_{i}^{\prime})}{len(\hat{Y}_{i}^{\prime\prime})}$
$CROUGE\text{-}L=\frac{1}{N}\sum_{i=1}^{N}\frac{2r_{i}^{\prime}r_{i}^{\prime\prime}}{r_{i}^{\prime}+r_{i}^{\prime\prime}},$
where $LCS(\cdot,\cdot)$ denotes the longest common subsequence of tokens between two LLM outputs, and $len(\hat{Y})$ denotes the number of tokens in an LLM output. The CROUGE-L metric effectively uses ROUGE-L to assess similarity as the longest common subsequence (LCS) relative to generated text length.
[…]
Counterfactual BLEU (CBLEU).
This work introduces CBLEU, defined as the average BLEU score (Papineni et al., 2002) over counterfactually generated output pairs. The full calculation of CBLEU is as follows:
$precision_{b}(\hat{Y}_{i}^{\prime},\hat{Y}_{i}^{\prime\prime})=\frac{\sum_{snt\in\hat{Y}_{i}^{\prime}}\sum_{b{\text{-}}gram\in snt}\min(C(b{\text{-}}gram,\hat{Y}_{i}^{\prime}|\hat{Y}_{i}^{\prime\prime}),C(b{\text{-}}gram,\hat{Y}_{i}^{\prime\prime}))}{\sum_{\tilde{snt}\in\hat{Y}_{i}^{\prime}}\sum_{b{\text{-}}gram\in\tilde{snt}}C(b{\text{-}}gram,\hat{Y}_{i}^{\prime})}$
$BLEU(\hat{Y}_{i}^{\prime},\hat{Y}_{i}^{\prime\prime})=\min(1,\exp\{1-\frac{len(\hat{Y}_{i}^{\prime\prime})}{len(\hat{Y}_{i}^{\prime})}\})(\prod_{b=1}^{4}precision_{b}(\hat{Y}_{i}^{\prime},\hat{Y}_{i}^{\prime\prime}))^{1/4}$
$CBLEU=\frac{1}{N}\sum_{i=1}^{N}\min(BLEU(\hat{Y}_{i}^{\prime},\hat{Y}_{i}^{\prime\prime}),BLEU(\hat{Y}_{i}^{\prime\prime},\hat{Y}_{i}^{\prime})),$
where $snt$ denotes a sentence in an LLM output, $len(\hat{Y})$ denotes the number of tokens in an LLM output, $C(b{\text{-}}gram,\hat{Y}_{i}^{\prime})$ denotes the number of times $b{\text{-}}gram$ appears in $\hat{Y}_{i}^{\prime}$ and $C(b{\text{-}}gram,\hat{Y}_{i}^{\prime}|\hat{Y}_{i}^{\prime\prime})$ denotes the number of times $b{\text{-}}gram$ appears in $\hat{Y}_{i}^{\prime}$ given that it also appears in $\hat{Y}_{i}^{\prime\prime}$ (Papineni et al., 2002; goo, ). For symmetry, the minimum of these two BLEU scores for each counterfactual pair is obtained before averaging. For the same reasons as with CROUGE-L, practitioners should mask protected attribute words in counterfactual output pairs before computing CBLEU.
Counterfactual Cosine Similarity (CCS).
Given a sentence transformer $\mathbf{V}:\mathcal{Y}\xrightarrow{}\mathbb{R}^{d}$, this work defines CCS as:
$CCS=\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{V}(Y_{i}^{\prime})\cdot\mathbf{V}(Y_{i}^{\prime\prime})}{\lVert\mathbf{V}(Y_{i}^{\prime})\rVert\lVert\mathbf{V}(Y_{i}^{\prime\prime})\rVert},$
i.e. the average cosine similarity (Singhal and Google, 2001) between counterfactually generated output pairs for an LLM use case.
3.1.3.2 Counterfactual Sentiment Bias
Counterfactual sentiment metrics measure the sentiment consistency across counterfactually generated pairs of output. To achieve this, these metrics leverage a pre-trained sentiment classifier $Sm:\mathcal{Y}\xrightarrow[]{}[0,1]$. This section outlines two counterfactual sentiment metrics: Strict Counterfactual Sentiment Parity, proposed by Huang et al. (2020), and an extension of this metric introduced in this work called Weak Counterfactual Sentiment Parity. [^19] Huang et al. (2020) use Google Cloud sentiment API and a BERT-based sentiment classifier. Both metrics have a range of values of $[0,1]$, with smaller values indicating a higher degree of fairness.
Strict Counterfactual Sentiment Parity (SCSP) (Huang et al., 2020).
SCSP calculates Wasserstein-1 distance (Jiang et al., 2019) between the output distributions of a sentiment classifier applied to counterfactually generated LLM outputs:
$\mathcal{W}_{1}=\mathbb{E}_{\tau\sim\mathcal{U}(0,1)}|P(Sm(\hat{Y}^{\prime})>\tau)-P(Sm(\hat{Y}^{\prime\prime})>\tau)|,$
where $\mathcal{U}(0,1)$ denotes the uniform distribution. Above, $\mathbb{E}_{\tau\sim\mathcal{U}(0,1)}$ is calculated empirically on a sample of counterfactual response pairs $(\hat{Y}_{1}^{\prime},\hat{Y}_{1}^{\prime\prime}),...,(\hat{Y}_{N}^{\prime},\hat{Y}_{N}^{\prime\prime})$ generated by $\mathcal{M}$, from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ drawn from $\mathcal{P}_{X|\mathcal{A}}$.
Weak Counterfactual Sentiment Parity (WCSP).
This study presents WCSP, defined as the discrepancy in predicted sentiment rates by a sentiment classifier when applied to counterfactually generated LLM outputs, given a threshold $\tau$ for binarizing sentiment scores.
$WCSP=\frac{1}{N}\sum_{i=1}^{N}|P(Sm(\hat{Y}_{i}^{\prime})>\tau)-P(Sm(\hat{Y}_{i}^{\prime\prime})>\tau)|.$
In practice, practitioners may select an appropriate value of $\tau$ depending on stakeholder values and the sentiment classifier being used.
3.2 Fairness Metrics for Classification Use Cases
It is well-established that classification models can produce unfair outcomes for certain protected attribute groups (Saleiro et al., 2018; Bellamy et al., 2018; Weerts et al., 2023; Feldman et al., 2014; Hardt et al., 2016; Mehrabi et al., 2019). Let a classification LLM use case be defined as an LLM tasked with classification, denoted as $\mathcal{M}^{(c)}$, and a population of prompts $\mathcal{P}_{X}$. Note that the metrics introduced in this section are confined to binary classification use cases, where $\mathcal{M}^{(c)}:\mathcal{X}\xrightarrow{}\{0,1\},$ with the understanding that evaluating fairness for multiclass classification is a straightforward extension from the binary case (Rouzot et al., 2023).
For the remainder of this section, assume each prompt in $\mathcal{P}_{X}$ for a given classification LLM use case corresponds to a protected attribute group. Under this assumption, traditional machine learning fairness metrics (Bellamy et al., 2018; Saleiro et al., 2018; Weerts et al., 2023) may be applied (Czarnowska et al., 2021). Accordingly, these metrics are defined on binary predictions $\hat{Y}_{1},...,\hat{Y}_{N}$, generated from a sample of prompts $X_{1},...,X_{N}\in\mathcal{P}_{X}$, with some metrics also incorporating corresponding ground truth values $Y_{1},...,Y_{N}$. These metrics effectively assess group fairness (see Section 2.2), with choice of statistical outcome measure $B$ depending on stakeholder values (e.g. the relative cost of false negatives vs. false positives).
This section distinguishes between representation fairness metrics, calculated using only predictions, and error-based fairness metrics, calculated using both predictions and ground truth values. The formulation of each fairness metric involves calculating the absolute difference between a pair of group-level metrics. This calculation yields a range of values between 0 and 1, where smaller values signify a higher level of fairness. [^20] Note that while this work considers group fairness in terms of differences between protected attribute groups, as in Gallegos et al. (2023); Bellamy et al. (2018), analogous metrics can be calculated as ratios (Saleiro et al., 2018).
3.2.1 Representation Fairness Metrics for Binary Classification
Representation fairness metrics aim to determine whether protected attribute groups are adequately represented in the positive predictions generated by classifier. It is recommended that practitioners reserve this set of metrics for classification LLM use cases for which group-level predicted prevalence rates, i.e. the proportion of predictions belonging to the positive class, should be approximately equal. [^21] For example, when predicting which applicants are qualified for a job it may be desirable to target equal rates of positive predictions for males and females. However, when predicting certain diseases, this may not be a desirable model behavior. A single representation fairness metric is defined below: Demographic Parity.
Demographic Parity (DP) (Dwork et al., 2011).
DP calculates the absolute difference in group-level predicted prevalence rates:
$DP=|{P(\hat{Y}=1|G=G^{\prime})}-{P(\hat{Y}=1|G=G^{\prime\prime})}|,$
where $\hat{Y}$ denotes a model prediction and $P(\cdot)$ denotes the empirical probability based on predictions generated from a sample prompts drawn from $\mathcal{P}_{X}$.
3.2.2 Error-Based Fairness Metrics for Binary Classification
Error-based fairness metrics aim to determine whether disparities in model performance exist across protected attribute groups. For error-based fairness, two metrics focused on false negatives, False Negative Rate Difference and False Omission Rate Difference, and two metrics focused on false positives, False Positive Rate Difference and False Discovery Rate Difference, are introduced. Following Saleiro et al. (2018), it is recommended that the choice between these two types of error-based fairness metrics be informed by whether the interventions assigned by the model are assistive (meaning false negatives are undesirable) or punitive (meaning false positives are undesirable) in nature.
False Negative Rate Difference (FNRD) (Bellamy et al., 2018).
FNRD measures the absolute difference in group-level false negative rates:
$FNRD=|{P(\hat{Y}=0|Y=1,G=G^{\prime})}-{P(\hat{Y}=0|Y=1,G=G^{\prime\prime})}|,$
where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ denotes the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$. Note that false negative rate measures the proportion of actual positives $(Y=1)$ that are falsely classified as negative $(\hat{Y}=0)$. FNRD is equivalent to the equal opportunity difference metric proposed by Hardt et al. (2016).
False Omission Rate Difference (FORD) (Bellamy et al., 2018).
FORD measures the absolute difference in group-level false omission rates:
$FORD=|{P(Y=1|\hat{Y}=0,G=G^{\prime})}-{P(Y=1|\hat{Y}=0,G=G^{\prime\prime})}|,$
where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ denotes the empirical probability based predictions generated from on a sample of prompts drawn from $\mathcal{P}_{X}$. Instead of concentrating on actual positives, false omission rate calculates the percentage of predicted negatives ($\hat{Y}=0$) that are misclassified. Thus, similar to the FNRD, a higher FORD indicates a greater difference in the likelihood of false negatives across groups.
False Positive Rate Difference (FPRD) (Bellamy et al., 2018).
FPRD measures the absolute difference in group-level false positive rates:
$FPRD=|{P(\hat{Y}=1|Y=0,G=G^{\prime})}-{P(\hat{Y}=1|Y=0,G=G^{\prime\prime})}|,$
where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ denotes the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$. Note that false positive rate measures the percentage of actual negatives ($Y=0$) being incorrectly predicted as positive $(\hat{Y}=1)$.
False Discovery Rate Difference (FDRD) (Bellamy et al., 2018).
FDRD measures the absolute difference in group-level false discovery rates:
$FDRD=|{P(Y=0|\hat{Y}=1,G=G^{\prime})}-{P(Y=0|\hat{Y}=1,G=G^{\prime\prime})}|,$
where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ denotes the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$. Rather than considering actual negatives, false discovery rate calculates the proportion of predicted positives ($\hat{Y}=1$) that are incorrectly classified. Hence, as with FPRD, a higher FDRD indicates a larger disparity in the likelihood of false positives across groups.
3.2.3 Multiclass Fairness Metrics
For multiclass classifiers, the framework proposed here follows the fairness guidelines provided by Rouzot et al. (2023) and hence recommends conducting class-wise fairness assessments using the appropriate binary classification fairness metrics, as per sections 3.2.1, 3.2.2, on each of the ‘sensitive’ classes. In particular, Rouzot et al. (2023) characterize sensitive classes as outcomes having significant impact on the lives of individuals to whom the model is applied.
3.3 Fairness Metrics for Recommendation Use Cases
Zhang et al. (2023) have shown that LLMs tasked with recommendation can exhibit discrimination when exposed to protected attribute information in the input prompts. Let a recommendation LLM use case be defined as an LLM tasked with recommendation, denoted as $\mathcal{M}^{(R)}$, and a population of prompts $\mathcal{P}_{X}$. Specifically, $\mathcal{M}^{(R)}:\mathcal{X}\xrightarrow{}\mathcal{R}^{K}$ maps a prompt $X\in\mathcal{X}$ to an ordered $K$-tuple $\hat{R}\in\mathcal{R}^{K}$ of distinct recommendations from a set of possible recommendations $\mathcal{R}$.
This section outlines a set of fairness metrics for recommendation LLM use cases, as proposed by Zhang et al. (2023). To ensure consistency with the metrics discussed in Section 3.1.3, this section presents modified versions of these metrics to be pairwise in nature, rather than attribute-wise. Given two protected attribute groups $G^{\prime},G^{\prime\prime},$ and an LLM use case $(\mathcal{M}^{(R)},\mathcal{P}_{X})$, these metrics assess similarity in counterfactually generated recommendation lists. Below, each metric is defined according to responses generated from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ which are drawn from $\mathcal{P}_{X|\mathcal{A}}$. In particular, three metrics are introduced: Jaccard Similarity, Search Result Page Misinformation Score at K, and Pairwise Ranking Accuracy Gap at K. Each of these metrics ranges in value from 0 to 1, with larger values indicating a greater degree of fairness.
Jaccard Similarity at K (Jaccard-K) (Zhang et al., 2023).
Below, a pairwise version of Jaccard-K is introduced. This metric calculates the average Jaccard Similarity (Han et al., 2011)—the ratio of the intersection cardinality to the union cardinality—among pairs of counterfactually generated recommendation lists. Formally, this metric is computed as follows:
$Jaccard\text{-}K=\frac{1}{N}\sum_{i=1}^{N}\frac{|\hat{R}_{i}^{\prime}\cap\hat{R}_{i}^{\prime\prime}|}{|\hat{R}_{i}^{\prime}\cup\hat{R}_{i}^{\prime\prime}|},$
where $\hat{R}_{i}^{\prime},\hat{R}_{i}^{\prime\prime}$ respectively denote the generated lists of recommendations by $\mathcal{M}(X;\theta)$ from the counterfactual input pair $(X_{i}^{\prime},X_{i}^{\prime\prime})$. Note that this metric does not account for ranking differences between the two lists (Zhang et al., 2023).
Search Result Page Misinformation Score at K (SERP-K) (Zhang et al., 2023).
Adapted from Tomlein et al. (2021), SERP-K reflects the similarity of two lists, considering both overlap and ranks. A modified version of SERP-K, adapted for pairwise application, is introduced as follows:
$\psi(X_{i}^{\prime},X_{i}^{\prime\prime})=\sum_{v\in\hat{R}_{i}^{\prime}}\frac{I(v\in\hat{R}_{i}^{\prime\prime})*(K-rank(v,\hat{R}_{i}^{\prime})+1)}{K*(K+1)/2},$
$SERP\text{-}K=\frac{1}{N}\sum_{i=1}^{N}\min(\psi(X_{i}^{\prime},X_{i}^{\prime\prime}),\psi(X_{i}^{\prime\prime},X_{i}^{\prime}))$
where $\hat{R}_{i}^{\prime},\hat{R}_{i}^{\prime\prime}$ respectively denote the generated lists of recommendations by $\mathcal{M}(X;\theta)$ from the counterfactual input pair $(X_{i}^{\prime},X_{i}^{\prime\prime})$, $v$ is a recommendation from $\hat{R}_{i}^{\prime}$, and $rank(v,\hat{R}_{i}^{\prime})$ denotes the rank of $v$ in $\hat{R}_{i}^{\prime}$. Note that the use of $\min(\cdot,\cdot)$ is included to achieve symmetry.
Pairwise Ranking Accuracy Gap at K (PRAG-K) (Zhang et al., 2023).
Adapted from Beutel et al. (2019), PRAG-K reflects the similarity in pairwise ranking between two recommendation results. A pairwise version of PRAG-K is presented as follows:
$rankmatch_{i}(v_{1},v_{2})=I(rank(v_{1},\hat{R}_{i}^{\prime})<rank(v_{2},\hat{R}_{i}^{\prime}))*I(rank(v_{1},\hat{R}_{i}^{\prime\prime})<rank(v_{2},\hat{R}_{i}^{\prime\prime}))$
$\eta(X_{i}^{\prime},X_{i}^{\prime\prime})=\sum_{v_{1},v_{2}\in\hat{R}_{i}^{\prime}\\ v_{1}\neq v_{2}}\frac{I(v_{1}\in\hat{R}_{i}^{\prime\prime})*rankmatch_{i}(v_{1},v_{2})}{K*(K+1)},$
$PRAG\text{-}K=\frac{1}{N}\sum_{i=1}^{N}\min(\eta(X_{i}^{\prime},X_{i}^{\prime\prime}),\eta(X_{i}^{\prime\prime},X_{i}^{\prime})),$
where $\hat{R}_{i}^{\prime},\hat{R}_{i}^{\prime\prime}$ respectively denote the generated lists of recommendations by $\mathcal{M}(X;\theta)$ from the counterfactual input pair $(X_{i}^{\prime},X_{i}^{\prime\prime})$, $v_{1},v_{2}$ are recommendations from $\hat{R}_{i}^{\prime}$, and $rank(v,\hat{R}_{i})$ denotes the rank of $v$ in $\hat{R}_{i}$. As with SERP-K, $\min(\cdot,\cdot)$ is used to achieve symmetry.
Table 2: Bias and Fairness Evaluation Metrics
Evaluation Metric Required Input
Toxicity
   Expected Maximum Toxicity 25 generations per prompt
   Toxicity Probability 25 generations per prompt
   Toxic Fraction 1 (or more) generation per prompt
Stereotype
   Stereotypical Associations 1 (or more) generation per prompt
   Co-occurrence Bias Score 1 (or more) generation per prompt
   Expected Maximum Stereotype 25 generations per prompt
   Stereotype Probability 25 generations per prompt
   Stereotype Fraction 1 (or more) generation per prompt
Counterfactual Fairness (Generated Text)
   Counterfactual ROUGE-L 1 (or more) counterfactual pair of generations per prompt
   Counterfactual BLEU 1 (or more) counterfactual pair of generations per prompt
   Counterfactual Cosine Similarity 1 (or more) counterfactual pair of generations per prompt
   Weak Counterfactual Sentiment Parity 1 (or more) counterfactual pair of generations per prompt
   Strict Counterfactual Sentiment Parity 1 (or more) counterfactual pair of generations per prompt
Allocational harms
   Demographic Parity Binary predictions and associated protected attribute groups
   False Negative Rate Difference Binary predictions, ground truth values, and associated protected attributed groups
   False Omission Rate Difference Binary predictions, ground truth values, and associated protected attributed groups
   False Positive Rate Difference Binary predictions, ground truth values, and associated protected attributed groups
   False Discovery Rate Difference Binary predictions, ground truth values, and associated protected attributed groups
Counterfactual Fairness (Recommendation)
   Jaccard-K Counterfactual pairs of generated recommendation lists of length $K$
   SERP-K Counterfactual pairs of generated recommendation lists of length $K$
   PRAG-K Counterfactual pairs of generated recommendation lists of length $K$
4 A Unified Framework for Bias and Fairness Assessments of LLM Use Cases
In general, bias and fairness assessments of LLM use cases do not require satisfying all possible evaluation metrics. Instead, practitioners should prioritize and concentrate on a relevant subset of metrics that align with their use case. To demystify metric choice for these assessments, this section introduces a decision framework that enables practitioners to determine suitable choices of bias and fairness evaluation metrics, drawing inspiration from the classification fairness framework proposed by Saleiro et al. (2018).
In defining the proposed decision framework for selecting LLM bias and fairness evaluation metrics, the scope is restricted to use cases for which prompts can be sampled from a known population and the task is well-defined. Use cases are categorized into three distinct groups based on task: 1) text generation and summarization, 2) classification, and 3) recommendation. The framework includes suitable metrics for each category to assess the potential bias and fairness risks that align with the specific characteristics of the use case. The mapping of these metrics is illustrated in Figure 1, and a thorough discussion of this mapping is provided below. For a comprehensive list of bias and fairness evaluation metrics, the reader may refer to Table 2.
First, consider text generation and summarization. For this collection of use cases, an important factor in determining relevant bias and fairness metrics is whether the use case upholds FTU, meaning that prompts do not include any mentions of protected attribute words. If FTU cannot be achieved for this type of use case, it is recommended that practitioners include counterfactual fairness and stereotype metrics, as respectively outlined in Sections 3.1.3 and 3.1.2, in their assessments. [^22] Note that counterfactual similarity may be too strict in some contexts, particularly in scenarios where textual content ought to differ across protected attribute groups (e.g. clinical contexts). For these scenarios, practitioners may opt to focus only on counterfactual sentiment and omit counterfactual similarity metrics from their assessment.  [^23] Note that stereotype risk, while low, may still exist even if FTU is satisfied. Hence, practitioners may wish to conduct stereotype assessments in these scenarios. Additionally, it is recommended that all text generation and summarization use cases undergo toxicity evaluation, as outlined in Section 3.1.1, regardless of whether or not FTU is achieved.
For classification use cases, a modified version of the decision framework proposed by Saleiro et al. (2018) is adopted. This framework can be applied to any classification use case where inputs correspond to protected attribute groups. Following Saleiro et al. (2018), the following approach is recommended: if fairness necessitates that model predictions exhibit approximately equal predicted prevalence across different groups, representation fairness metrics should be used; Otherwise, error-based fairness metrics should be used. For error-based fairness, practitioners should focus on disparities in false negatives (positives), assessed by FNRD and FORD (FPRD and FDRD), if the model is used to assign assistive (punitive) interventions. [^24] In the context of fairness, if interventions are punitive, and hence can hurt individuals, it is undesirable for a model to produce false positives disproportionately for any protected attribute group. Analogously, having a model produce false negatives disproportionately for any protected attribute group is undesirable in the case of assistive interventions. If inputs cannot be mapped to a protected attribute, meaning they are not person-level inputs and they satisfy FTU, then a fairness assessment is not applicable.
Third, for recommendation use cases, counterfactual unfairness is a risk if FTU cannot be satisfied, as shown by Zhang et al. (2023). Note that counterfactual invariance may not be a desirable property for certain recommendation use cases. For instance, it may be preferred to recommend different products for male vs. female customers. Hence, if counterfactual invariance is a desired property, it is recommended that recommendation use cases not satisfying FTU be assessed for counterfactual unfairness in recommendations using the metrics outlined in 3.3. On the other hand, if a recommendation use case satisfies FTU or if counterfactual invariance is not desired, then a fairness assessment is not applicable.
It is important to note that the framework presented here is intended for use cases in which the volume of LLM responses makes exhaustive human review infeasible. Accordingly, it is important to note that in situations where the practitioner manually reviews each generated output, the proposed evaluations may be unnecessary if the concerns regarding bias and fairness can be effectively addressed by the output reviewer themselves.
Figure 1: Bias and Fairness Evaluation Framework for LLM Use Cases
5 Conclusion
This paper proposes an actionable decision framework for selecting bias and fairness evaluation metrics for LLM use cases, introducing several new evaluation metrics as part of the framework. This work addresses two gaps in the current literature. First, to the best of the author’s knowledge, the current literature does not offer a framework for selecting bias and fairness evaluation metrics for LLM use cases. To fill this gap, the proposed framework draws inspiration from Saleiro et al. (2018) and incorporates use case characteristics and stakeholder values to guide the selection of evaluation metrics. Second, this framework tackles limitations of existing LLM bias and fairness evaluation approaches that rely on benchmark data sets containing predefined prompts. Instead, the approach outlined in this work involves using actual prompts from the practitioner’s use case. By considering both prompt-risk and the assigned task of the LLM, this approach provides a more customized risk assessment for the practitioner’s specific use case. Furthermore, the proposed framework aims to enhance practicality and ease of implementation, as all evaluation metrics are computed solely from the LLM output.
Acknowledgements
Thank you to Mohit Singh Chauhan, Blake Aber, Piero Ferrante, Xue (Crystal) Gu, Almira Pillay, Zeya Ahmad, and Vasistha Singhal Vinod for your helpful suggestions.

### v2

@@
 1 Introduction
-The versatility of current Large Language Models (LLMs) in handling various tasks (Minaee et al., 2024; Liu et al., 2023; Ray, 2023) presents challenges when it comes to evaluating bias and fairness at the model level.
-Existing approaches primarily focus on assessing risk using benchmark data sets containing predefined prompts (Gehman et al., 2020; Dhamala et al., 2021; Nozza et al., 2021; Smith et al., 2022; Parrish et al., 2021; Li et al., 2020), masked tokens (Zhao et al., 2018; Rudinger et al., 2018; Nadeem et al., 2021; Levy et al., 2021), or unmasked sentences (Nangia et al., 2020; Barikeri et al., 2021; Jiao et al., 2023; Felkner et al., 2023), assuming that these adequately capture specific bias or fairness risks (Gallegos et al., 2023).
-However, these assessments are likely to overestimate the risk for use cases where the population of prompts is low risk (Wang et al., 2024).
+The versatility of current Large Language Models (LLMs) in handling various tasks (Minaee et al., 2024; Liu et al., 2023; Ray, 2023) presents challenges when it comes to evaluating bias and fairness at the model level (Anthis et al., 2024).
+Existing approaches primarily focus on assessing risk using benchmark data sets containing predefined prompts (Gehman et al., 2020; Dhamala et al., 2021; Nozza et al., 2021; Smith et al., 2022; Parrish et al., 2021; Li et al., 2020; Wang et al., 2024a), masked tokens (Zhao et al., 2018; Rudinger et al., 2018; Nadeem et al., 2021; Levy et al., 2021), or unmasked sentences (Nangia et al., 2020; Barikeri et al., 2021; Jiao et al., 2023; Felkner et al., 2023), assuming that these adequately capture specific bias or fairness risks (Gallegos et al., 2023).
+However, these assessments are likely to overestimate the risk for use cases where the population of prompts is low risk (Wang et al., 2024b).
 Moreover, to the best of the author’s knowledge, the current literature does not provide a framework for effectively aligning LLM use cases with suitable metrics for evaluating bias and fairness.
@@
 (2018), the framework proposed in this work enables practitioners to map an LLM use case to an appropriate set of bias and fairness evaluation metrics by considering relevant characteristics of the use case and stakeholder values.
-This evaluation approach is unique in that it incorporates actual prompts from the practitioner’s use case, taking into account the prompt-specific risks that have been demonstrated to significantly increase the likelihood of biased and unfair outcomes (Wang et al., 2024).
+This evaluation approach is unique in that it incorporates actual prompts from the practitioner’s use case, taking into account the prompt-specific risks that have been demonstrated to significantly increase the likelihood of biased and unfair outcomes (Wang et al., 2024b).
 By constraining the scope to focused use cases, where prompts are derived from a known population and the task is well-defined, this framework is specifically designed to customize the risk assessment for a specific application.
-To introduce the framework, this study first provides formal definitions of bias and fairness desiderata for LLMs from the literature and segment these definitions by risk category.
+To introduce the framework, this study first provides formal definitions of bias and fairness desiderata for LLMs from the literature and segments these definitions by risk category.
 Subsequently, those risks are mapped to a taxonomy of use cases focused on large-scale applications, where human-in-the-loop may be infeasible.
@@
 The discussions in this paper explore concepts of bias and fairness in relation to an arbitrary ‘protected attribute’, encompassing examples such as sex, race, age, and sexual orientation, among others.
-These concepts are established formally by using mathematical notation whenever possible and strives to prioritize consistency in notation throughout, even if it may differ from the notation used in the papers cited.
+Formal establishment of these concepts is achieved through the use of mathematical notation whenever feasible.
+The aim is to maintain consistency in notation throughout, even if it deviates from the notation used in the referenced papers.
 2.1 Preliminary Definitions
@@
 2.2.3 Counterfactual Fairness
-In many contexts, it is undesirable for an LLM to generate substantially different output as a result of different protected attribute words contained in the input prompts, all else equal (Huang et al., 2020; Nozza et al., 2021; Wang et al., 2024).
+In many contexts, it is undesirable for an LLM to generate substantially different output as a result of different protected attribute words contained in the input prompts, all else equal (Huang et al., 2020; Nozza et al., 2021; Wang et al., 2024b).
 Following previous work (Huang et al., 2020; Garg et al., 2019), this concept is hereafter referred to as (lack of) counterfactual fairness.
-Depending on context and stakeholder values, the practitioner may wish to assess an LLM for differences in overall content or sentiment resulting from inclusion of different protected attribute words in a prompt.
+Depending on context and stakeholder values, the practitioner may wish to assess an LLM use case for differences in overall content or sentiment resulting from inclusion of different protected attribute words in a prompt.
 Below, this section provides a formal definition of the corresponding fairness desiderata known as counterfactual invariance, adapted from Gallegos et al.
@@
 Accounting for prompt-specific risk is important to accurately reflect the risk of a specific use case, as Wang et al.
-(2024) find in their evaluations that toxicity probability is 26 to 101 times higher when dealing with toxic prompts compared to non-toxic prompts.
+(2024b) find in their evaluations that toxicity probability is 26 to 101 times higher when dealing with toxic prompts compared to non-toxic prompts.
 Moreover, Goldfarb-Tarrant et al.
@@
 While Gallegos et al.
-(2023) further segment generated-text metrics into three subcategories (namely distribution metrics, classifier metrics, and lexicon metrics), this work segments these metrics according to the risk taxonomy outlined in 2.2.
+(2023) further segment generated-text metrics into three subcategories (namely distribution metrics, classifier metrics, and lexicon metrics), this work segments these metrics according to the risk taxonomy outlined in Section 2.2.
 In particular, generated-text bias and fairness metrics are segmented into the following categories: toxicity metrics, stereotype metrics, and counterfactual fairness metrics.
-Toxicity metrics leverage a pre-trained toxicity classifier, such as Perspective API [^12] https://perspectiveapi.com, to assign a toxicity score to an LLM’s output (Chowdhery et al., 2022; Lees et al., 2022; Wang et al., 2024; Bommasani et al., 2023; Gehman et al., 2020).
+Toxicity metrics leverage a pre-trained toxicity classifier, such as Perspective API [^12] https://perspectiveapi.com, to assign a toxicity score to an LLM’s output (Chowdhery et al., 2022; Lees et al., 2022; Wang et al., 2024b; Bommasani et al., 2023; Gehman et al., 2020).
 Stereotype metrics assess the relative co-occurrence of stereotype words with protected attribute words (Bordia and Bowman, 2019; Bommasani et al., 2023) or leverage a pre-trained stereotype classifier to assign a stereotype score to an LLM’s output (Zekun et al., 2023).
-Lastly, counterfactual fairness metrics assess fairness through a causal lens by exploiting variation in mentions of protected attributes in otherwise-identical prompts (Huang et al., 2020).
+Lastly, counterfactual fairness metrics assess fairness through a causal lens by exploiting variation in mentions of protected attribute groups in otherwise-identical prompts (Huang et al., 2020).
 Select metrics are detailed in each category below.
@@
 Specifically, these metrics are evaluated on a set of $m$ responses per prompt, $\{\hat{Y}_{ij}\}_{j=1}^{m}$, generated by $\mathcal{M}$, from a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$.
-In practice, variation in responses for the same prompt can be achieved by choosing a non-zero temperature for the LLM or by utilizing various random seeds (Wang et al., 2024).
+In practice, variation in responses for the same prompt can be achieved by choosing a non-zero temperature for the LLM or by utilizing various random seeds (Wang et al., 2024b).
 Below, formal definitions of three toxicity metrics are provided: Expected Maximum Toxicity, Toxicity Probability, and Toxic Fraction.
@@
 $EMT=\frac{1}{N}\sum_{i=1}^{N}\max_{1\leq j\leq m}T(\hat{Y}_{ij}).$
-Note that the standard choice of $m$ for this metric is $m=25$ (Gehman et al., 2020; Wang et al., 2024).
+Note that the standard choice of $m$ for this metric is $m=25$ (Gehman et al., 2020; Wang et al., 2024b).
 Toxicity Probability (TP) (Gehman et al., 2020).
@@
 $TP=\frac{1}{N}\sum_{i=1}^{N}I(\max_{1\leq j\leq m}T(\hat{Y}_{ij})\geq 0.5).$
-As with EMT, the standard choice of $m$ for this metric is $m=25$ (Gehman et al., 2020; Wang et al., 2024).
+As with EMT, the standard choice of $m$ for this metric is $m=25$ (Gehman et al., 2020; Wang et al., 2024b).
 Toxicity Fraction (TF) (Bommasani et al., 2023).
@@
 Co-Occurence Bias Score (COBS) (Bordia and Bowman, 2019).
-Given two protected attribute groups $G^{\prime},G^{\prime\prime}$ with associated sets of protected attribute words $A^{\prime},A^{\prime\prime}$, a set of stereotypical words $W$, and an LLM use case $(\mathcal{M},\mathcal{P}_{X})$, the full calculation of COBS is as follows:
-$P(w|A)=\frac{\sum_{i=1}^{N}cooccur(w,A|\hat{Y}_{i})/\sum_{i=1}^{N}\sum_{\tilde{w}\in\hat{Y}_{i}}cooccur(\tilde{w},A|\tilde{Y}_{i})}{\sum_{i=1}^{N}\sum_{a\in A}C(a,\hat{Y}_{i})/\sum_{i=1}^{N}\sum_{\tilde{w}\in\tilde{Y}_{i}}C(\tilde{w},\hat{Y}_{i})}$
-$COBS=\frac{1}{|W|}\sum_{w\in W}\log\frac{P(w|A^{\prime})}{P(w|A^{\prime\prime})}.$
-Above, $C(x,\hat{Y}_{i})$ denotes the count of $x$ in $\hat{Y}_{i}$ and $\tilde{Y}_{i}$ represents the LLM output $\hat{Y}_{i}$, with words from $A^{\prime}\cup A^{\prime\prime}$ and stop words excluded.
-The co-occurrence function $cooccur(w,A|\hat{Y})$ computes a weighted count of words from $A$ that are found within a context window centered around $w$, each time $w$ appears in $\hat{Y}$.
-[^15] To specify the co-occurrence function $cooccur(w,A|\hat{Y})$, Bordia and Bowman (2019) define a fixed or infinite context window centered at a target word $w$.
-In the case of fixed context (Bordia and Bowman (2019) recommend including 10 words before and 10 words after), each word in the context window receives equal weight.
-In the case of an infinite context window, each word receives weight $\beta^{dist},$ where $dist$ denotes the number of tokens between the two tokens of interest (Bordia and Bowman (2019) use $\beta=0.95$).
-In both cases, only words from the context window included in $A$ are included in the weighted count.
-In words, COBS computes the relative likelihood that an LLM $\mathcal{M}$ generates output having co-occurrence of $w\in W$ with $A^{\prime}$ versus $A^{\prime\prime}$.
-This metric has a range of possible values of [,][$\infty$,$\infty$][ - ∞ , ∞ ], with values closer to 0 signifying a greater degree of fairness.
+Given two protected attribute groups $G^{\prime},G^{\prime\prime}$ with associated sets of protected attribute words $A^{\prime},A^{\prime\prime}$, a set of stereotypical words $W$, a set of stop words $\mathcal{S}$, and an LLM use case $(\mathcal{M},\mathcal{P}_{X})$, the full calculation of COBS is as follows:
+$cooccur(w,A|\hat{Y})=\sum_{w_{j},w_{k}\in\hat{Y},w_{j}\neq w_{k}}I(w_{j}=w)\cdot I(w_{k}\in A)\cdot\beta^{dist(w_{j},w_{k})}$
+$P(w|A)=\frac{\sum_{i=1}^{N}cooccur(w,A|\hat{Y}_{i})/\sum_{i=1}^{N}\sum_{\tilde{w}\in\hat{Y}_{i}}cooccur(\tilde{w},A|\hat{Y}_{i})\cdot I(\tilde{w}\notin\mathcal{S}\cup\mathcal{A})}{\sum_{i=1}^{N}\sum_{a\in A}C(a,\hat{Y}_{i})/\sum_{i=1}^{N}\sum_{\tilde{w}\in\hat{Y}_{i}}C(\tilde{w},\hat{Y}_{i})\cdot I(\tilde{w}\notin\mathcal{S}\cup\mathcal{A})}$
+$COBS=\frac{1}{|W|}\sum_{w\in W}\log\frac{P(w|A^{\prime})}{P(w|A^{\prime\prime})},$
+where $C(x,\hat{Y}_{i})$ denotes the count of $x$ in $\hat{Y}_{i}$ and $dist(w_{j},w_{k})$ denotes the number of tokens between $w_{j}$ and $w_{k}$.
+Above, the co-occurrence function $cooccur(w,A|\hat{Y})$ computes a weighted count of words from $A$ that are found within a context window centered around $w$, each time $w$ appears in $\hat{Y}$.
+[^15] Although Bordia and Bowman (2019) introduce two versions of this metric—one with a fixed-context window and another with an infinite context window—only the version with the infinite context window is incorporated into this framework.
+In their work, Bordia and Bowman (2019) use $\beta=0.95.$ Note that the functions $cooccur(\tilde{w},A|\hat{Y}_{i})$ and $C(\tilde{w},\hat{Y}_{i})$ are multiplied by zero for $\tilde{w}\in\mathcal{S}\cup\mathcal{A}$ in order to exclude stop words and protected attribute words from these counts.
+Put simply, COBS computes the relative likelihood that an LLM $\mathcal{M}$ generates output having co-occurrence of $w\in W$ with $A^{\prime}$ versus $A^{\prime\prime}$.
+This metric has a range of possible values of (,)($\infty$,$\infty$)( - ∞ , ∞ ), with values closer to 0 signifying a greater degree of fairness.
 Stereotypical Associations (SA) (Bommasani et al., 2023).
-Consider a set of protected attribute groups $\mathcal{G}$, an associated set of protected attribute lexicons $\mathcal{A}$, and an associated set of stereotypical words $W$.
-Additionally, let $C(x,\hat{Y})$ denote the number of times that the word $x$ appears in the output $\hat{Y}$, $I(\cdot)$ denote the indicator function, $P^{\text{ref}}$ denote a reference distribution, and $TVD$ denote total variation difference.
+Consider a set of protected attribute groups $\mathcal{G}$, an associated set of protected attribute group lexicons $\mathcal{A}$, and an associated set of stereotypical words $W$.
+Additionally, let $C(x,\hat{Y})$ denote the number of times that the word $x$ appears in the output $\hat{Y}$, $P^{\text{ref}}$ denote a reference distribution, and $TVD$ denote total variation difference.
 [^16] The reference distribution recommended by Bommasani et al.
@@
 Counterfactual similarity metrics measure the similarity in outputs generated from counterfactual input pairs according to a specified invariance metric $\upsilon$, i.e.
-$\upsilon(\mathcal{M}(X^{\prime};\theta),\mathcal{M}(X^{\prime\prime};\theta)).$ These metrics effectively assess whether the LLM use case satisfies the counterfactual invariance property defined in 2.2.
+$\upsilon(\mathcal{M}(X^{\prime};\theta),\mathcal{M}(X^{\prime\prime};\theta)).$ These metrics effectively assess whether the LLM use case satisfies the counterfactual invariance property defined in Section 2.2.
 One such example of $\upsilon$ is exact match (Rajpurkar et al., 2016), but (Gallegos et al., 2023) argue that this metric is too strict.
@@
 SCSP calculates Wasserstein-1 distance (Jiang et al., 2019) between the output distributions of a sentiment classifier applied to counterfactually generated LLM outputs:
-$\mathcal{W}_{1}=\mathbb{E}_{\tau\sim\mathcal{U}(0,1)}|P(Sm(\hat{Y}^{\prime})>\tau)-P(Sm(\hat{Y}^{\prime\prime})>\tau)|,$
+$SCSP=\mathbb{E}_{\tau\sim\mathcal{U}(0,1)}|P(Sm(\hat{Y}^{\prime})>\tau)-P(Sm(\hat{Y}^{\prime\prime})>\tau)|,$
 where $\mathcal{U}(0,1)$ denotes the uniform distribution.
@@
 Weak Counterfactual Sentiment Parity (WCSP).
-This study presents WCSP, defined as the discrepancy in predicted sentiment rates by a sentiment classifier when applied to counterfactually generated LLM outputs, given a threshold $\tau$ for binarizing sentiment scores.
-$WCSP=\frac{1}{N}\sum_{i=1}^{N}|P(Sm(\hat{Y}_{i}^{\prime})>\tau)-P(Sm(\hat{Y}_{i}^{\prime\prime})>\tau)|.$
+This study presents WCSP, defined as the difference in predicted sentiment rates by a sentiment classifier applied to counterfactually generated LLM output pairs.
+Given a threshold $\tau$ for binarizing sentiment scores, this metric is introduced as follows:
+$WCSP=\left\lvert\frac{1}{N}\sum_{i=1}^{N}I(Sm(\hat{Y}_{i}^{\prime})>\tau)-\frac{1}{N}\sum_{i=1}^{N}I(Sm(\hat{Y}_{i}^{\prime\prime})>\tau)\right\rvert.$
 In practice, practitioners may select an appropriate value of $\tau$ depending on stakeholder values and the sentiment classifier being used.
@@
 Accordingly, these metrics are defined on binary predictions $\hat{Y}_{1},...,\hat{Y}_{N}$, generated from a sample of prompts $X_{1},...,X_{N}\in\mathcal{P}_{X}$, with some metrics also incorporating corresponding ground truth values $Y_{1},...,Y_{N}$.
-These metrics effectively assess group fairness (see Section 2.2), with choice of statistical outcome measure $B$ depending on stakeholder values (e.g.
+These metrics effectively assess group fairness (see Section 2.2) between two protected attribute groups $G^{\prime}$ and $G^{\prime\prime}$, with choice of statistical outcome measure $B$ depending on stakeholder values (e.g.
 the relative cost of false negatives vs.
@@
 However, when predicting certain diseases, this may not be a desirable model behavior.
-A single representation fairness metric is defined below: Demographic Parity.
+A single representation fairness metric, known as Demographic Parity, is defined below.
 Demographic Parity (DP) (Dwork et al., 2011).
@@
 $DP=|{P(\hat{Y}=1|G=G^{\prime})}-{P(\hat{Y}=1|G=G^{\prime\prime})}|,$
-where $\hat{Y}$ denotes a model prediction and $P(\cdot)$ denotes the empirical probability based on predictions generated from a sample prompts drawn from $\mathcal{P}_{X}$.
+where $\hat{Y}$ denotes a model prediction and $P(\cdot)$ represents the empirical probability based on predictions generated from a sample prompts drawn from $\mathcal{P}_{X}$.
 3.2.2 Error-Based Fairness Metrics for Binary Classification
 Error-based fairness metrics aim to determine whether disparities in model performance exist across protected attribute groups.
-For error-based fairness, two metrics focused on false negatives, False Negative Rate Difference and False Omission Rate Difference, and two metrics focused on false positives, False Positive Rate Difference and False Discovery Rate Difference, are introduced.
+For error-based fairness, two metrics focused on false negatives, False Negative Rate Difference and False Omission Rate Difference, and two metrics focused on false positives, False Positive Rate Difference and False Discovery Rate Difference, are presented.
 Following Saleiro et al.
@@
 $FNRD=|{P(\hat{Y}=0|Y=1,G=G^{\prime})}-{P(\hat{Y}=0|Y=1,G=G^{\prime\prime})}|,$
-where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ denotes the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$.
+where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ represents the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$.
 Note that false negative rate measures the proportion of actual positives $(Y=1)$ that are falsely classified as negative $(\hat{Y}=0)$.
@@
 $FORD=|{P(Y=1|\hat{Y}=0,G=G^{\prime})}-{P(Y=1|\hat{Y}=0,G=G^{\prime\prime})}|,$
-where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ denotes the empirical probability based predictions generated from on a sample of prompts drawn from $\mathcal{P}_{X}$.
+where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ represents the empirical probability based predictions generated from on a sample of prompts drawn from $\mathcal{P}_{X}$.
 Instead of concentrating on actual positives, false omission rate calculates the percentage of predicted negatives ($\hat{Y}=0$) that are misclassified.
@@
 $FPRD=|{P(\hat{Y}=1|Y=0,G=G^{\prime})}-{P(\hat{Y}=1|Y=0,G=G^{\prime\prime})}|,$
-where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ denotes the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$.
+where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ represents the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$.
 Note that false positive rate measures the percentage of actual negatives ($Y=0$) being incorrectly predicted as positive $(\hat{Y}=1)$.
@@
 $FDRD=|{P(Y=0|\hat{Y}=1,G=G^{\prime})}-{P(Y=0|\hat{Y}=1,G=G^{\prime\prime})}|,$
-where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ denotes the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$.
+where $Y$ denotes ground truth value corresponding to $\hat{Y}$ and $P(\cdot)$ represents the empirical probability based on predictions generated from a sample of prompts drawn from $\mathcal{P}_{X}$.
 Rather than considering actual negatives, false discovery rate calculates the proportion of predicted positives ($\hat{Y}=1$) that are incorrectly classified.
@@
 Below, each metric is defined according to responses generated from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ which are drawn from $\mathcal{P}_{X|\mathcal{A}}$.
-In particular, three metrics are introduced: Jaccard Similarity, Search Result Page Misinformation Score at K, and Pairwise Ranking Accuracy Gap at K.
+In particular, three metrics are presented: Jaccard Similarity, Search Result Page Misinformation Score at K, and Pairwise Ranking Accuracy Gap at K.
 Each of these metrics ranges in value from 0 to 1, with larger values indicating a greater degree of fairness.
@@
 (2018).
-In defining the proposed decision framework for selecting LLM bias and fairness evaluation metrics, the scope is restricted to use cases for which prompts can be sampled from a known population and the task is well-defined.
+The scope of this framework is restricted to use cases for which prompts can be sampled from a known population and the task is well-defined.
 Use cases are categorized into three distinct groups based on task: 1) text generation and summarization, 2) classification, and 3) recommendation.
@@
 female customers.
-Hence, if counterfactual invariance is a desired property, it is recommended that recommendation use cases not satisfying FTU be assessed for counterfactual unfairness in recommendations using the metrics outlined in 3.3.
+Hence, if counterfactual invariance is a desired property, it is recommended that recommendation use cases not satisfying FTU be assessed for counterfactual unfairness in recommendations using the metrics outlined in Section 3.3.
 On the other hand, if a recommendation use case satisfies FTU or if counterfactual invariance is not desired, then a fairness assessment is not applicable.

### v3

@@
 Such biases can create or exacerbate unfair outcomes for certain groups within a protected attribute, including, but not limited to sex, race, sexual orientation, or age.
-This paper aims to provide a technical guide for practitioners to assess bias and fairness risks in LLM use cases.
-The main contribution of this work is a decision framework that allows practitioners to determine which metrics to use for a specific LLM use case.
-To achieve this, this study categorizes LLM bias and fairness risks, maps those risks to a taxonomy of LLM use cases, and then formally defines various metrics to assess each type of risk.
-As part of this work, several new bias and fairness metrics are introduced, including innovative counterfactual metrics as well as metrics based on stereotype classifiers.
-Instead of focusing solely on the model itself, the sensitivity of both prompt-risk and model-risk are taken into account by defining evaluations at the level of an LLM use case, characterized by a model and a population of prompts.
-Furthermore, because all of the evaluation metrics are calculated solely using the LLM output, the proposed framework is highly practical and easily actionable for practitioners.
-[^1] Note: The examples provided in this paper are purely hypothetical and are not intended to reflect the specific work or practices of the author-affiliated company.
+In this paper, we propose a decision framework that allows practitioners to determine which bias and fairness metrics to use for a specific LLM use case.
+To establish the framework, we define bias and fairness risks for LLMs, map those risks to a taxonomy of LLM use cases, and then define various metrics to assess each type of risk.
+Instead of focusing solely on the model itself, we account for both prompt-specific- and model-specific-risk by defining evaluations at the level of an LLM use case, characterized by a model and a population of prompts.
+Furthermore, because all of the evaluation metrics are calculated solely using the LLM output, our proposed framework is highly practical and easily actionable for practitioners.
+For streamlined implementation, all evaluation metrics included in the framework are offered in this paper’s companion Python toolkit, LangFair.
+Finally, our experiments demonstrate substantial variation in bias and fairness across use cases, underscoring the importance of use-case-level assessments.
+[^1] Note: The examples provided in this paper are purely hypothetical and are not intended to reflect the specific work, practices, or opinions of the author-affiliated company.
 They are used solely for illustrative purposes.
 Any resemblance to actual practices or projects is coincidental.
+All opinions are the author’s own.
 1 Introduction
-The versatility of current Large Language Models (LLMs) in handling various tasks (Minaee et al., 2024; Liu et al., 2023; Ray, 2023) presents challenges when it comes to evaluating bias and fairness at the model level (Anthis et al., 2024).
-Existing approaches primarily focus on assessing risk using benchmark data sets containing predefined prompts (Gehman et al., 2020; Dhamala et al., 2021; Nozza et al., 2021; Smith et al., 2022; Parrish et al., 2021; Li et al., 2020; Wang et al., 2024a), masked tokens (Zhao et al., 2018; Rudinger et al., 2018; Nadeem et al., 2021; Levy et al., 2021), or unmasked sentences (Nangia et al., 2020; Barikeri et al., 2021; Jiao et al., 2023; Felkner et al., 2023), assuming that these adequately capture specific bias or fairness risks (Gallegos et al., 2023).
-However, these assessments are likely to overestimate the risk for use cases where the population of prompts is low risk (Wang et al., 2024b).
-Moreover, to the best of the author’s knowledge, the current literature does not provide a framework for effectively aligning LLM use cases with suitable metrics for evaluating bias and fairness.
-This work aims to address these limitations by developing an actionable LLM bias and fairness evaluation framework defined at the use case level.
+The versatility of current Large Language Models (LLMs) in handling various tasks (Minaee et al., 2024; Liu et al., 2023; Ray, 2023) makes evaluating bias and fairness at the model level difficult (Anthis et al., 2024).
+Existing approaches largely rely on benchmark datasets containing predefined prompts (Gehman et al., 2020; Dhamala et al., 2021; Nozza et al., 2021; Smith et al., 2022; Parrish et al., 2021; Li et al., 2020; Wang et al., 2024a), masked tokens (Zhao et al., 2018; Rudinger et al., 2018; Nadeem et al., 2021; Levy et al., 2021), or unmasked sentences (Nangia et al., 2020; Barikeri et al., 2021; Jiao et al., 2023; Felkner et al., 2023), assuming that these adequately capture specific bias or fairness risks (Gallegos et al., 2023).
+However, these assessments do not account for prompt-specific risks that have been shown to significantly influence the likelihood of biased and unfair LLM responses (Wang et al., 2024b).
+[^2] Wang et al.
+(2024b) find in their evaluations that toxicity probability is 26 to 101 times higher when dealing with toxic prompts compared to non-toxic prompts.
+Moreover, to the best of our knowledge, the current literature does not provide a framework for effectively aligning LLM use cases with suitable metrics for evaluating bias and fairness.
+To address these limitations, we propose an LLM bias and fairness evaluation framework defined at the use case level.
 Drawing inspiration from the classification fairness framework proposed by Saleiro et al.
-(2018), the framework proposed in this work enables practitioners to map an LLM use case to an appropriate set of bias and fairness evaluation metrics by considering relevant characteristics of the use case and stakeholder values.
-This evaluation approach is unique in that it incorporates actual prompts from the practitioner’s use case, taking into account the prompt-specific risks that have been demonstrated to significantly increase the likelihood of biased and unfair outcomes (Wang et al., 2024b).
-By constraining the scope to focused use cases, where prompts are derived from a known population and the task is well-defined, this framework is specifically designed to customize the risk assessment for a specific application.
-To introduce the framework, this study first provides formal definitions of bias and fairness desiderata for LLMs from the literature and segments these definitions by risk category.
-Subsequently, those risks are mapped to a taxonomy of use cases focused on large-scale applications, where human-in-the-loop may be infeasible.
-Lastly, for each risk category, various bias and fairness evaluation metrics are detailed, with discussions provided on their input requirements, calculation methods, the risks they assess, and circumstances under which they should be applied.
-As part of this work, a variety of novel bias and fairness metrics are introduced.
-This includes innovative counterfactual adaptations of recall-oriented understudy for gisting evaluation (ROUGE) (Lin, 2004), bilingual evaluation understudy (BLEU) (Papineni et al., 2002), and cosine similarity (Singhal and Google, 2001), as well a set of stereotype classifier-based metrics that are adapted from analogous toxicity classifier-based metrics.
-For practical reasons, this study limits the selection of LLM bias and fairness metrics to those requiring only LLM generated output as inputs.
-This includes 1) generated text metrics, which take a generated set of tokens as input (Gallegos et al., 2023), 2) recommendation fairness metrics, calculated on a set of LLM-provided recommendations (Zhang et al., 2023), and 3) classification fairness metrics, which are already well-established in the machine learning fairness literature (Bellamy et al., 2018; Saleiro et al., 2018; Weerts et al., 2023; Hardt et al., 2016; Feldman et al., 2014; Mehrabi et al., 2019).
-Due to practical limitations of the input requirements, the framework omits both embedding-based metrics, which are computed using an LLM’s hidden vector representations of words or sentences (Islam et al., 2016; May et al., 2019; Guo and Caliskan, 2020), and probability-based metrics, which leverage predicted token probabilities from an LLM (Webster et al., 2020; Kurita et al., 2019; Ahn and Oh, 2021; Kaneko and Bollegala, 2021; Salazar et al., 2019; Barikeri et al., 2021; Nangia et al., 2020; Nadeem et al., 2021).
-To further support the rationale behind the selection of evaluation metrics, it is important to note that metrics focused on the downstream task, consistent with the metrics incorporated in this framework, have been shown to be more reliable than metrics derived from embeddings or token probabilities (Goldfarb-Tarrant et al., 2020; Delobelle et al., 2021).
-The remainder of this paper is organized as follows.
-Section 2 formally defines various notions of bias and fairness in LLMs.
-Section 3 discusses various techniques for conducting bias and fairness assessments in LLMs.
-Section 4 offers a framework for choosing among LLM bias and fairness metrics based on use case characteristics and stakeholder values.
-Finally, Section 5 offers concluding remarks.
+(2018), our framework enables practitioners to map an LLM use case to an appropriate set of bias and fairness evaluation metrics by considering the task, relevant characteristics of the prompts, and stakeholder values.
+This evaluation approach is unique in that it follows a bring-your-own-prompts approach, in which metrics are computed from LLM responses to actual prompts from the practitioner’s use case.
+Our framework is designed for well-defined use cases, where prompts are sampled from a known population, allowing for bias and fairness assessments that are customized for a specific application.
+To introduce the framework, we first define bias and fairness risks for LLMs from the literature and map those risks to a taxonomy of use cases.
+For each risk category, we then present various evaluation metrics and discuss their input requirements, computation methods, the risks they assess, and circumstances under which they should be applied.
+As part of this work, we also introduce a variety of novel bias and fairness metrics.
+Specifically, these new metrics include counterfactual adaptations of recall-oriented understudy for gisting evaluation (ROUGE) (Lin, 2004), bilingual evaluation understudy (BLEU) (Papineni et al., 2002), and cosine similarity (Singhal and Google, 2001), as well as a set of stereotype classifier-based metrics that are adapted from analogous toxicity classifier-based metrics.
+For practical reasons, we limit the selection of LLM bias and fairness metrics to those requiring only LLM generated output for computation.
+To streamline implementation of the framework, all included bias and fairness metrics are offered by this paper’s companion Python toolkit, LangFair.
+[^3] https://github.com/cvs-health/langfair In practice, users provide a sample of prompts from their use case and their LLM of choice, and LangFair streamlines generation of LLM responses and computes applicable metrics for their use case.
+This toolkit offers a model-agnostic, user-friendly way to implement our evaluation framework for real-world use cases.
+Lastly, we conduct a series of experiments to evaluate bias and fairness in several text generation and summarization use cases.
+In particular, we construct 6 unique use cases, characterized by three sets of prompts and two LLMs.
+We find substantial variation in bias and fairness across use cases, underscoring the importance of use-case-level assessments.
 2 Bias and Fairness Risks for LLM Use Cases
-This section formally defines several pre-requisite terms and concepts upon which subsequent sections rely, several of which are adapted from those provided by Gallegos et al.
+In this section, we define various terms upon which subsequent sections rely, several of which are adapted from those provided by Gallegos et al.
 (2023).
@@
 (2023).
-The approach outlined below adopts modified versions of these definitions to better fit the tone and purpose.
-For further reading on biases in LLMs, the reader may refer to (Gallegos et al., 2023; Blodgett et al., 2020; Kumar et al., 2023; Li et al., 2024; Chu et al., 2024; Ferrara, 2023; Ranaldi et al., 2023; Kotek et al., 2023; Wu and Aji, 2023; Li et al., 2023; Ray, 2023; Nozza et al., 2022; Zhuo et al., 2023).
-The discussions in this paper explore concepts of bias and fairness in relation to an arbitrary ‘protected attribute’, encompassing examples such as sex, race, age, and sexual orientation, among others.
-Formal establishment of these concepts is achieved through the use of mathematical notation whenever feasible.
-The aim is to maintain consistency in notation throughout, even if it deviates from the notation used in the referenced papers.
+We adopt modified versions of these definitions to better fit the tone and purpose.
+For further reading on biases in LLMs, the reader may refer to (Gallegos et al., 2023; Blodgett et al., 2020; Kumar et al., 2023; Li et al., 2024; Chu et al., 2024; Ferrara, 2023; Ranaldi et al., 2023; Kotek et al., 2023; Wu and Aji, 2023; Li et al., 2023a; Ray, 2023; Nozza et al., 2022; Zhuo et al., 2023).
+Note that, throughout this paper, we explore concepts of bias and fairness in relation to an arbitrary ‘protected attribute’, encompassing examples such as sex, race, age, and sexual orientation, among others.
 2.1 Preliminary Definitions
-This section provides preliminary definitions to be used throughout the subsequent sections.
+Below we provide several preliminary definitions to be used throughout the subsequent sections.
 Large Language Model (LLM).
 An LLM $\mathcal{M}:\mathcal{X}\xrightarrow{}\mathcal{Y}$ is a pre-trained, transformer-based model that maps a text sequence $X\in\mathcal{X}$ to an output $\hat{Y}\in\mathcal{Y}$, where $\mathcal{X}$ denotes the set of all possible text inputs (i.e.
-prompts) and the form of $\hat{Y}$ is specific to the LLM and the use case (Gallegos et al., 2023).
+user prompts) and the form of $\hat{Y}$ is specific to the LLM and the use case (Gallegos et al., 2023).
 [^5] Following Gallegos et al.
@@
 Population of Prompts.
-A population of prompts, denoted $\mathcal{P}_{X}$, is a collection of LLM inputs.
-To characterize well-defined use cases, subsequent sections refer to a ‘known population of prompts’, indicating that practitioners possess information about the prompt domain and are able draw representative samples from $\mathcal{P}_{X}$.
+A population of prompts, denoted $\mathcal{P}_{X}$, is a collection of LLM inputs (user prompts).
+To characterize well-defined use cases, we subsequently refer to a ‘known population of prompts’, indicating that practitioners possess information about the prompt domain and are able draw representative samples from $\mathcal{P}_{X}$.
 For instance, a population of prompts might consist of clinical notes, where each individual prompt includes a collection of notes, accompanied by specific instructions for the LLM to generate a summary (Chuang et al., 2024).
@@
 In the interest of concise notation, LLM use cases will be hereafter denoted as $(\mathcal{M},\mathcal{P}_{X}).$ An LLM use case is evaluated on a finite set of responses generated by $\mathcal{M}(X;\theta)$ from a sample of $N$ prompts $X_{1},...,X_{N}$, drawn from the population $\mathcal{P}_{X}$.
-Protected Attribute Groups.
+Protected Attribute Group.
 A protected attribute group $G\in\mathcal{G}$ represents a subset of people characterized by a shared identity trait, where $\mathcal{G}$ is a partition (Gallegos et al., 2023).
@@
 A protected attribute group lexicon $A\in\mathcal{A}$ is a collection of words that correspond to protected attribute group $G\in\mathcal{G}$ [^7] Hereafter, in a slight abuse of notation, $\mathcal{G}$ (and analogously $\mathcal{A}$) can denote either a partition of or the union of each $G\in\mathcal{G}$ (and analogously each $A\in\mathcal{A}$).
-[^6] One example, provided by Bommasani et al.
+[^8] One example, provided by Liang et al.
 (2023), of protected attribute words for the protected attribute group ‘males’ is { ‘he’, ‘son’, ‘his’, ‘him’, ‘father’, ‘man’, ‘boy’, ‘himself’, ‘male’, ‘brother’, ‘sons’, ‘fathers’, ‘men’, ‘boys’, ‘males’, ‘brothers’, ‘uncle’, ‘uncles’, ‘nephew’, ‘nephews’ }.
 [^9] The examples that consider male vs.
-female are used in accordance with the convention established by previous studies such as Bordia and Bowman (2019) and Bommasani et al.
+female are used in accordance with the convention established by previous studies such as Bordia and Bowman (2019) and Liang et al.
 (2023).
 It is important to note that this usage is solely for the purpose of consistency and does not intend to imply that there are only two genders.
-The acknowledgement of a broader spectrum of gender identities is recognized and respected.
+The acknowledgment of a broader spectrum of gender identities is recognized and respected.
 Counterfactual Input Pair.
@@
 2.2 LLM Bias and Fairness Risks
-This section presents formal definitions of various notions of bias and fairness applicable to LLMs.
-A taxonomy for these definitions is proposed, organized by risk category.
-Specifically, the risk categories included in this taxonomy are toxicity, stereotyping, counterfactual fairness, and allocational harms.
+In this section, we define various bias and fairness risks applicable to LLMs and define corresponding desiderata.
+Namely, these risks include toxicity, stereotyping, counterfactual fairness, and allocational harms.
 2.2.1 Toxicity
-Drawing from the definitions of toxicity and derogatory language outlined in the survey conducted by Gallegos et al.
-(2023), let the definition of toxic text encompass any offensive language that 1) launches attacks, issues threats, or incites hate or violence against a social group, or 2) includes the usage of pejorative slurs, insults, or any other forms of expression that specifically target and belittle a social group.
-To formalize this, a corresponding fairness desideratum is introduced below.
+Following Gallegos et al.
+(2023), we define toxic text as any offensive language that 1) launches attacks, issues threats, or incites hate or violence against a social group, or 2) includes the usage of pejorative slurs, insults, or any other forms of expression that specifically target and belittle a social group.
+To formalize this, we introduce a corresponding desideratum known as non-toxicity.
 Non-Toxicity.
@@
 2.2.2 Stereotyping
-Stereotyping is an important type of social bias that should be considered in the context of LLMs (Bommasani et al., 2023; Bordia and Bowman, 2019; Zekun et al., 2023).
-This study follows the work by Gallegos et al.
-(2023), in which stereotypes are defined as negative generalizations about a protected attribute group, often reflected by differences in frequency with which various groups are linked to stereotyped terms (Bommasani et al., 2023).
-To formalize this notion, two fairness desiderata are considered, which are also proposed by Gallegos et al.
-(2023): equal group associations (EGA) and equal neutral associations (ENA).
+Stereotyping is an important type of social bias that should be considered in the context of LLMs (Liang et al., 2023; Bordia and Bowman, 2019; Zekun et al., 2023).
+We follow Gallegos et al.
+(2023) and define stereotypes as negative generalizations about a protected attribute group, often reflected by differences in frequency with which various groups are linked to stereotyped terms (Liang et al., 2023).
+The corresponding desideratum, proposed by Gallegos et al.
+(2023), is known as equal group associations.
 Equal Group Associations (Gallegos et al., 2023).
 For two protected attribute groups $G^{\prime},G^{\prime\prime},$ and a set of neutral words $W$, an LLM use case $(\mathcal{M},\mathcal{P}_{X})$ satisfies equal group associations if, for each $w\in W,P(w\in\hat{Y}|\hat{Y}\cap A^{\prime}\neq\emptyset)=P(w\in\hat{Y}|\hat{Y}\cap A^{\prime\prime}\neq\emptyset)$.
-Put simply, equal group associations requires that each neutral word in $W$ is equally likely to be contained in the output of $\mathcal{M}$ regardless of which protected attribute group is mentioned.
-Equal Neutral Associations (Gallegos et al., 2023).
-For two protected attribute groups $G^{\prime},G^{\prime\prime},$ with respective associated lexicons $A^{\prime},A^{\prime\prime},$ and a set of neutral words $W$, an LLM use case $(\mathcal{M},\mathcal{P}_{X})$ satisfies Equal Neutral Associations if, $P(A^{\prime}\cap\hat{Y}\neq\emptyset|\hat{Y}\cap W\neq\emptyset)=P(A^{\prime\prime}\cap\hat{Y}\neq\emptyset|\hat{Y}\cap W\neq\emptyset).$ In other words, Equal Neutral Associations requires that co-occurrence of protected attribute words with a set of neutral words $W$ is equally probable for both protected attribute groups.
-[^9] In practice, the set $W$ referenced in the definitions of equal group associations and equal neutral associations typically contains frequently stereotyped terms such as professions or adjectives (Bordia and Bowman, 2019; Bommasani et al., 2023).
+Put simply, equal group associations requires that each neutral word in $W$ is equally likely to be contained in the output of $\mathcal{M}$, regardless of which protected attribute group is mentioned.
 2.2.3 Counterfactual Fairness
 In many contexts, it is undesirable for an LLM to generate substantially different output as a result of different protected attribute words contained in the input prompts, all else equal (Huang et al., 2020; Nozza et al., 2021; Wang et al., 2024b).
-Following previous work (Huang et al., 2020; Garg et al., 2019), this concept is hereafter referred to as (lack of) counterfactual fairness.
+Following previous work (Huang et al., 2020; Garg et al., 2019), we refer to this concept as (lack of) counterfactual fairness.
 Depending on context and stakeholder values, the practitioner may wish to assess an LLM use case for differences in overall content or sentiment resulting from inclusion of different protected attribute words in a prompt.
-Below, this section provides a formal definition of the corresponding fairness desiderata known as counterfactual invariance, adapted from Gallegos et al.
+Below, we present the corresponding fairness desideratum, known as counterfactual invariance, adapted from Gallegos et al.
 (2023).
@@
 (2023).
+Note that some metrics that assess counterfactual invariance assess similarity, rather than dissimilarity, meaning the inequality will be reversed.
 2.2.4 Allocational Harms
@@
 (2023) define as an unequal distribution of resources or opportunities among different protected attribute groups, have been widely studied in the machine learning fairness literature (Saleiro et al., 2018; Bellamy et al., 2018; Weerts et al., 2023; Kamishima et al., 2012; Zhang et al., 2018; Hardt et al., 2016; Feldman et al., 2014; Pleiss et al., 2017; Kamiran et al., 2012; Agarwal et al., 2018; Kamiran and Calders, 2011; Chouldechova, 2016).
-The corresponding fairness desideratum, is known as group fairness, defined formally below.
+In this work, we measure allocational harms based on group fairness, defined below.
 Group Fairness.
@@
 2.3 Mapping Bias and Fairness Risks to LLM Use Cases
-Section 2.1 establishes the characterization of a well-defined LLM use case based on a model and a known population of prompts.
-[^11] Note that use case segmentation proposed in this work is by task, which can be controlled through various means such as fine-tuning of the LLM, providing examples with few-shot prompting, or incorporating instructions in system or user prompts.
-To set up the decision framework, this section segments use cases by task, according to the following three categories: 1) text generation and summarization, 2) classification, and 3) recommendation.
+In section 2.1, we characterize an LLM use case based on a model and a known population of prompts.
+Here, we categorize use cases into three task-based groups: 1) text generation and summarization, 2) classification, and 3) recommendation.
+[^12] Note that our case segmentation is by task, which can be controlled through various means such as fine-tuning of the LLM, providing examples with few-shot prompting, or incorporating instructions in system or user prompts.
+[^13] Note that our proposed framework is intended for large-scale applications, in which the volume of generated responses makes exhaustive human review impractical.
+For scenarios in which the practitioner manually evaluates each generated output, the evaluations we propose may be unnecessary if concerns related to bias and fairness can be effectively addressed by the individual who is reviewing the outputs.
 Descriptions and examples are provided in Table 1.
-The bias and fairness evaluation framework proposed in this work is intended for large-scale applications, in which the volume of generated responses makes exhaustive human review impractical.
-It is important to note that, for scenarios in which the practitioner manually evaluates each generated output, the evaluations proposed here may be unnecessary if concerns related to bias and fairness can be effectively addressed by the individual who is reviewing the outputs.
-Table 1: Taxonomy of LLM Use Cases and Associated Bias/Fairness Risks
-Use Case Category Description Examples Bias/Fairness Risk
-Text Generation and Summarization LLM generates text outputs that are not constrained to a predefined set of classes or list elements Create personalized outreach messages to individuals; Summarize clinical notes Toxic text, stereotypes*, counterfactual fairness*
-Classification LLM classifies a text input among a pre-defined set of classes Classify intent of customer support inquiries to assign assistance; Classify customer feedback as positive or negative to assign follow-ups Allocational harms**
-Recommendation LLM generates lists of recommendations Generate lists of recommended products; Generate lists of recommended news articles Counterfactual fairness*
-*Risk is applicable if FTU is not satisfied. Counterfactual fairness may not be relevant in certain contexts.
-**Risk is applicable if text inputs correspond to a protected attribute.
-Text Generation and Summarization.
-First, consider use cases where an LLM generates text outputs that are not constrained to a predefined set of classes (e.g., positive vs.
+2.3.1 Text Generation and Summarization
+We first consider use cases where an LLM generates text outputs that are not constrained to a predefined set of classes (e.g., positive vs.
 negative) or list elements (e.g., products to recommend).
-For the sake of brevity, this group of use cases will be hereafter referred to as “text generation and summarization," acknowledging that this category can encompass additional use cases including, but not limited to, machine translation, retrieval augmented generation (RAG), and question-answering.
-An example of a text generation use case could be utilizing an LLM to compose personalized messages for customer outreach.
-Similarly, a summarization use case may involve employing an LLM to extract pertinent information or provide summaries from clinical notes.
-These use cases carry the potential risk of generating toxic text in their outputs.
-Moreover, if these use cases fail to meet the criteria of FTU, meaning that the prompts include references to protected attributes, they also pose the risk of perpetuating stereotypes or exhibiting counterfactual unfairness.
-Classification.
+For the sake of brevity, we will hereafter refer to this group of use cases as “text generation and summarization," acknowledging that this category can encompass additional tasks including machine translation, question-answering, and others.
+One example of this type of use case is using an LLM to compose personalized messages for customer outreach.
+Use cases in this category carry the risk of generating toxic text in their outputs.
+Moreover, if these use cases fail to satisfy FTU, meaning that the prompts include mentions of protected attributes, they also pose the risk of perpetuating stereotypes or exhibiting counterfactual unfairness.
+2.3.2 Classification
 LLMs have been widely used for text classification (Sun et al., 2023; Widmann and Wich, 2022; Bonikowski et al., 2022; Howard and Ruder, 2018; Sun et al., 2019; Chai et al., 2020; Chen et al., 2020; Lin et al., 2021).
 In the context of bias and fairness, it is important to distinguish whether the text inputs can be mapped to a protected attribute, either by containing direct mentions of a protected attribute group, or in the case of person-level prompt granularity, corresponding to individuals belonging to certain protected attribute groups.
-For instance, one example of a person-level classification use case could involve utilizing an LLM to classify customer feedback as positive or negative in order to assign appropriate follow-ups.
-Similar to traditional person-level classification challenges in machine learning, these use cases present the risk of allocational harms.
+For instance, using an LLM to classify customer feedback as positive or negative in order to assign appropriate follow-ups would be an example of a person-level classification use case.
+Similar to traditional person-level classification problems in machine learning, these use cases present the risk of allocational harms.
 On the other hand, classification use cases that do not involve person-level data and satisfy FTU are not subject to these bias and fairness risks.
-Recommendation.
+2.3.3 Recommendation
 Recommendation is another potential application of LLMs (Bao et al., 2023; Gao et al., 2023), such as using an LLM to recommend products to customers.
@@
 (2023) show that LLMs used as recommendation engines can discriminate when exposed to protected attribute information.
-Given this concern, it follows that LLM recommendation use cases pose the risk of counterfactual unfairness if they do not satisfy FTU.
+It follows that LLM recommendation use cases pose the risk of counterfactual unfairness if they do not satisfy FTU.
 3 Bias and Fairness Evaluation Metrics
-This section introduces a range of evaluation metrics segmented by the use case task and its applicable bias and fairness risks.
-The proposed framework encompasses three distinct use case categories: 1) text generation and summarization, 2) classification, and 3) recommendation.
-For each category, bias and fairness evaluation metrics are included that address the applicable risks.
-Accounting for prompt-specific risk is important to accurately reflect the risk of a specific use case, as Wang et al.
-(2024b) find in their evaluations that toxicity probability is 26 to 101 times higher when dealing with toxic prompts compared to non-toxic prompts.
-Moreover, Goldfarb-Tarrant et al.
-(2020); Delobelle et al.
-(2021) have demonstrated that evaluation metrics which take into account the LLM’s task provide a more accurate reflection of the associated risk compared to metrics based on embeddings or token probabilities.
-Accordingly, the objective of this work is to assess the risks associated with a particular use case, taking into consideration not only the LLM used but also the task at hand and the prompt population.
-Following the characterization of a use case based on a model and a known population of prompts, as established in Section 2.1, each metric definition presented below is contextualized within an evaluation sample of size $N$ drawn from a known population of prompts $\mathcal{P}_{X}$.
-3.1 Bias Metrics for Text Generation and Summarization Use Cases
-Gallegos et al.
-(2023) propose a detailed taxonomy of evaluation metrics for bias evaluations in LLMs.
-They partition these metrics into three main categories: embedding-based metrics, probability-based metrics, and generated-text metrics.
-While embedding- and probability-based metrics require access to an LLM’s upstream architecture, generated-text metrics instead treat an LLM like a black box and can be easily calculated from LLM output alone.
-Hence, given their ease of application, only generated-text metrics are included in the framework proposed here.
-While Gallegos et al.
-(2023) further segment generated-text metrics into three subcategories (namely distribution metrics, classifier metrics, and lexicon metrics), this work segments these metrics according to the risk taxonomy outlined in Section 2.2.
-In particular, generated-text bias and fairness metrics are segmented into the following categories: toxicity metrics, stereotype metrics, and counterfactual fairness metrics.
-Toxicity metrics leverage a pre-trained toxicity classifier, such as Perspective API [^12] https://perspectiveapi.com, to assign a toxicity score to an LLM’s output (Chowdhery et al., 2022; Lees et al., 2022; Wang et al., 2024b; Bommasani et al., 2023; Gehman et al., 2020).
-Stereotype metrics assess the relative co-occurrence of stereotype words with protected attribute words (Bordia and Bowman, 2019; Bommasani et al., 2023) or leverage a pre-trained stereotype classifier to assign a stereotype score to an LLM’s output (Zekun et al., 2023).
+Our proposed framework encompasses three distinct use case categories: 1) text generation and summarization, 2) classification, and 3) recommendation.
+For each category, we present various evaluation metrics that address the applicable bias and fairness risks.
+For practical reasons, we limit the selection of metrics to those requiring only LLM generated output for computation.
+[^14] Due to practical limitations of the input requirements, we omit from the framework both embedding-based metrics, which are computed using an LLM’s hidden vector representations of words or sentences (Islam et al., 2016; May et al., 2019; Guo and Caliskan, 2020), and probability-based metrics, which leverage predicted token probabilities from an LLM (Webster et al., 2020; Kurita et al., 2019; Ahn and Oh, 2021; Kaneko and Bollegala, 2021; Salazar et al., 2019; Barikeri et al., 2021; Nangia et al., 2020; Nadeem et al., 2021).
+While these metrics require access to an LLM’s upstream architecture, the metrics we include in our framework do not.
+Importantly, we note that metrics focused on the downstream task, consistent with the metrics incorporated in this framework, have been shown to be more reliable than metrics derived from embeddings or token probabilities (Goldfarb-Tarrant et al., 2020; Delobelle et al., 2021).
+To ensure our metric definitions accurately reflect use-case-specific nature of our framework, we contextualize each metric within an evaluation sample of size $N$ drawn from a known population of prompts $\mathcal{P}_{X}$.
+3.1 Metrics for Text Generation and Summarization Use Cases
+We segment metrics for text generation and summarization use cases based on the applicable bias and fairness risks, as outlined in Section 2.2.
+Namely, this includes toxicity metrics, stereotype metrics, and counterfactual fairness metrics.
+Toxicity metrics leverage a pre-trained toxicity classifier, such as Perspective API [^15] https://perspectiveapi.com, to assign a toxicity score to an LLM’s output (Chowdhery et al., 2022; Lees et al., 2022; Wang et al., 2024b; Liang et al., 2023; Gehman et al., 2020).
+Stereotype metrics assess the relative co-occurrence of stereotype words with protected attribute words (Bordia and Bowman, 2019; Liang et al., 2023) or leverage a pre-trained stereotype classifier to assign a stereotype score to an LLM’s output (Zekun et al., 2023).
 Lastly, counterfactual fairness metrics assess fairness through a causal lens by exploiting variation in mentions of protected attribute groups in otherwise-identical prompts (Huang et al., 2020).
 Select metrics are detailed in each category below.
-[^13] Note that the discussion of generated text bias metrics in this paper is not exhaustive.
+[^16] We acknowledge that our set of bias and fairness metrics is not exhaustive.
 Rather, the intention is to select an appropriate set of evaluation metrics to address each risk to be assessed.
-Notable metrics omitted from this framework include HONEST (Nozza et al., 2021), score parity (Sicilia and Alikhani, 2023), Full Gen Bias (Smith et al., 2022), regard score (Sheng et al., 2019), demographic representations (Bommasani et al., 2023), psycholinguistic norms (Dhamala et al., 2021), and gender polarity (Dhamala et al., 2021).
+Notable metrics omitted from this framework include HONEST (Nozza et al., 2021), score parity (Sicilia and Alikhani, 2023), Full Gen Bias (Smith et al., 2022), regard score (Sheng et al., 2019), demographic representations (Liang et al., 2023), psycholinguistic norms (Dhamala et al., 2021), and gender polarity (Dhamala et al., 2021).
 3.1.1 Toxicity Metrics
-Toxicity metrics leverage a pre-trained toxicity classifier to assess general toxicity that is not specific to a particular protected attribute.
-Given a toxicity classifier $T:\mathcal{Y}\xrightarrow[]{}[0,1]$, this section defines toxicity metrics in the context of an LLM use case $(\mathcal{M},\mathcal{P}_{X})$.
-Specifically, these metrics are evaluated on a set of $m$ responses per prompt, $\{\hat{Y}_{ij}\}_{j=1}^{m}$, generated by $\mathcal{M}$, from a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$.
+Toxicity metrics leverage a pre-trained toxicity classifier, $T:\mathcal{Y}\xrightarrow[]{}[0,1]$, that maps a text input to a toxicity score.
+Here, we define toxicity metrics in the context of an LLM use case $(\mathcal{M},\mathcal{P}_{X})$.
+Specifically, these metrics are computed on a set of $m$ responses per prompt, $\{\hat{Y}_{ij}\}_{j=1}^{m}$, generated by $\mathcal{M}$, from a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$.
 In practice, variation in responses for the same prompt can be achieved by choosing a non-zero temperature for the LLM or by utilizing various random seeds (Wang et al., 2024b).
-Below, formal definitions of three toxicity metrics are provided: Expected Maximum Toxicity, Toxicity Probability, and Toxic Fraction.
-All three metrics range in values from 0 to 1, with smaller values indicating a greater degree of fairness.
+Below, we define three toxicity metrics: Expected Maximum Toxicity, Toxicity Probability, and Toxic Fraction.
+All three metrics range in value from 0 to 1, with smaller values indicating a greater degree of fairness.
 Expected Maximum Toxicity (EMT) (Gehman et al., 2020).
@@
 As with EMT, the standard choice of $m$ for this metric is $m=25$ (Gehman et al., 2020; Wang et al., 2024b).
-Toxicity Fraction (TF) (Bommasani et al., 2023).
+Toxicity Fraction (TF) (Liang et al., 2023).
 TF is defined as the fraction of generations that are classified as toxic:
 $TF=\frac{1}{N}\frac{1}{m}\sum_{i=1}^{N}\sum_{j=1}^{m}I(T(\hat{Y}_{ij})\geq 0.5),$
-This metric effectively estimates the likelihood that responses generated by $\mathcal{M}$ on prompts from $\mathcal{P}_{X}$ contain toxic text (Bommasani et al., 2023).
-Note that while the standard choice of $m$ for this metric is $m=1$ (Bommasani et al., 2023), a larger value of $m$ may be preferred in practice if sampling a large $N$ is infeasible.
+This metric effectively estimates the likelihood that responses generated by $\mathcal{M}$ on prompts from $\mathcal{P}_{X}$ contain toxic text (Liang et al., 2023).
+Note that while the standard choice of $m$ for this metric is $m=1$ (Liang et al., 2023), a larger value of $m$ may be preferred in practice if sampling a large $N$ is infeasible.
 3.1.2 Stereotype Metrics
@@
 [^17] Note that stereotype risk, while low, may still exist even if FTU is satisfied for an LLM use case.
-Among stereotype metrics, this work distinguishes between metrics based on co-occurrence of protected attribute words and stereotypical words, and metrics that leverage a stereotype classifier.
+Among stereotype metrics, we distinguish between metrics based on co-occurrence of protected attribute words and stereotypical words, and metrics that leverage a stereotype classifier.
 3.1.2.1 Co-occurrence-Based Metrics
-This section outlines a set of metrics that assess stereotype risk based on relative co-occurrence of protected attribute words with neutral words of interest.
-In particular, formal definitions of two co-occurrence-based stereotype metrics are provided: Co-Occurrence Bias Score and Stereotypical Associations.
+In this section, we outline a set of metrics that assess stereotype risk based on relative co-occurrence of protected attribute words with stereotype words of interest.
+These metrics effectively assess the degree to which equal group associations, as defined in Section 2, is satisfied.
+We define two co-occurrence-based stereotype metrics: Co-Occurrence Bias Score and Stereotypical Associations.
 Co-Occurence Bias Score (COBS) (Bordia and Bowman, 2019).
@@
 This metric has a range of possible values of (,)($\infty$,$\infty$)( - ∞ , ∞ ), with values closer to 0 signifying a greater degree of fairness.
-Stereotypical Associations (SA) (Bommasani et al., 2023).
+Stereotypical Associations (SA) (Liang et al., 2023).
 Consider a set of protected attribute groups $\mathcal{G}$, an associated set of protected attribute group lexicons $\mathcal{A}$, and an associated set of stereotypical words $W$.
 Additionally, let $C(x,\hat{Y})$ denote the number of times that the word $x$ appears in the output $\hat{Y}$, $P^{\text{ref}}$ denote a reference distribution, and $TVD$ denote total variation difference.
-[^16] The reference distribution recommended by Bommasani et al.
+[^19] The reference distribution recommended by Liang et al.
 (2023) is the uniform distribution.
@@
 In words, SA measures the relative co-occurrence of a set of stereotypically associated words across protected attribute groups.
+[^20] Note that while COBS and SA both assess equal group associations, COBS is computed pairwise, while SA is computed attribute-wise.
 SA ranges in value from 0 to 1, where smaller values indicate greater fairness.
@@
 [^21] Zekun et al.
-(2023) refer to https://huggingface.co/Narrativa/distilroberta-finetuned-stereotype-detection.
-This section introduces three new metrics by extending the toxicity metrics outlined in Section 3.1.1, leveraging a pre-trained stereotype classifier $St:\mathcal{Y}\xrightarrow[]{}[0,1]$ rather than a toxicity classifier.
+(2023) use https://huggingface.co/Narrativa/distilroberta-finetuned-stereotype-detection for their classifier.
+We introduce three stereotype metrics by extending the toxicity metrics outlined in Section 3.1.1, leveraging a pre-trained stereotype classifier, $St:\mathcal{Y}\xrightarrow[]{}[0,1]$, rather than a toxicity classifier.
 Namely, these metrics include: Expected Maximum Stereotype, Stereotype Probability, and Stereotype Fraction.
-All three metrics range in values from 0 to 1, with smaller values indicating a greater degree of fairness.
+All three metrics range in value from 0 to 1, with smaller values indicating a greater degree of fairness.
+We define these metrics below.
 Expected Maximum Stereotype (EMS).
@@
 effectively estimating the likelihood that responses generated by $\mathcal{M}$ on prompts from $\mathcal{P}_{X}$ contain stereotypes.
-Note that while the standard choice of $m$ for the analogous toxicity metric, EMT, is $m=1$ (Bommasani et al., 2023), a larger value of $m$ may be preferred in practice if sampling a large $N$ is infeasible.
+Note that while the standard choice of $m$ for the analogous toxicity metric, EMT, is $m=1$ (Liang et al., 2023), a larger value of $m$ may be preferred in practice if sampling a large $N$ is infeasible.
 3.1.3 Counterfactual Fairness Metrics
 Counterfactual metrics aim to assess differences in LLM output when different protected attributes are mentioned in input prompts, all else equal.
-Given two protected attribute groups $G^{\prime},G^{\prime\prime}$, these metrics are defined in the context of an LLM use case $(\mathcal{M},\mathcal{P}_{X})$.
+Given two protected attribute groups $G^{\prime},G^{\prime\prime}$, we define these metrics in the context of an LLM use case $(\mathcal{M},\mathcal{P}_{X})$.
 In particular, these metrics are evaluated on a sample of counterfactual response pairs $(\hat{Y}_{1}^{\prime},\hat{Y}_{1}^{\prime\prime}),...,(\hat{Y}_{N}^{\prime},\hat{Y}_{N}^{\prime\prime})$ generated by $\mathcal{M}$, from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ drawn from $\mathcal{P}_{X|\mathcal{A}}$.
@@
 It is important to note that the mapping does not need to be exhaustive, as long as it has sufficient coverage to generate a large sample of counterfactual input pairs.
-Note that, in scenarios where a large $N$ is infeasible, practitioners may opt to generate multiple response pairs per counterfactual input pair.
-These metrics, which are categorized into counterfactual similarity metrics and counterfactual sentiment metrics, respectively quantify the differences in text similarity and sentiment by leveraging the variations in LLM output observed across counterfactual input pairs.
+Note that, in scenarios where a large $N$ is infeasible, practitioners may opt to generate multiple response pairs per counterfactual input pair, as is done for toxicity metrics in Section 3.1.1.
+These metrics, which we categorize into counterfactual similarity metrics and counterfactual sentiment metrics, respectively quantify the differences in text similarity and sentiment by leveraging the variations in LLM output observed across counterfactual input pairs.
 Due to their reliance on mentions of protected attributes in input prompts, if FTU is satisfied for an LLM use case, these metrics need not be used.
@@
 $\upsilon(\mathcal{M}(X^{\prime};\theta),\mathcal{M}(X^{\prime\prime};\theta)).$ These metrics effectively assess whether the LLM use case satisfies the counterfactual invariance property defined in Section 2.2.
-One such example of $\upsilon$ is exact match (Rajpurkar et al., 2016), but (Gallegos et al., 2023) argue that this metric is too strict.
-This work introduces three, less stringent, counterfactual similarity metrics: Counterfactual ROUGE-L, Counterfactual BLEU, and Counterfactual Cosine Similarity, which are extensions of state-of-the-art text similarity metrics (Minaee et al., 2024; Lin, 2004; Papineni et al., 2002; Singhal and Google, 2001; Gomaa and Fahmy, 2013).
-While the first two assess similarity using token-sequence overlap, the the third assesses similarity using sentence embeddings.
-All three metrics range in values from 0 to 1, with larger values indicating a greater degree of fairness.
+One such example of $\upsilon$ is exact match (Rajpurkar et al., 2016), but Gallegos et al.
+(2023) argue that this metric is too strict.
+We introduce three, less stringent, counterfactual similarity metrics: Counterfactual ROUGE-L, Counterfactual BLEU, and Counterfactual Cosine Similarity, which are extensions of state-of-the-art text similarity metrics (Minaee et al., 2024; Lin, 2004; Papineni et al., 2002; Singhal and Google, 2001; Gomaa and Fahmy, 2013).
+The first two assess similarity using token-sequence overlap and range in value from 0 to 1.
+The third assesses similarity using sentence embeddings and ranges in value from -1 to 1.
+For each, larger values indicate a greater degree of fairness.
 Counterfactual ROUGE-L (CROUGE-L).
-This work introduces CROUGE-L, defined as the average ROUGE-L score (Lin, 2004) over counterfactually generated output pairs. The full calculation of CROUGE-L is as follows:
+We introduce CROUGE-L, defined as the average ROUGE-L score (Lin, 2004) over counterfactually generated output pairs. The full calculation of CROUGE-L is as follows:
 $r_{i}^{\prime}=\frac{LCS(\hat{Y}_{i}^{\prime},\hat{Y}_{i}^{\prime\prime})}{len(\hat{Y}_{i}^{\prime})}$
@@
 Counterfactual BLEU (CBLEU).
-This work introduces CBLEU, defined as the average BLEU score (Papineni et al., 2002) over counterfactually generated output pairs. The full calculation of CBLEU is as follows:
+We define CBLEU as the average BLEU score (Papineni et al., 2002) over counterfactually generated output pairs. The full calculation of CBLEU is as follows:
 $precision_{b}(\hat{Y}_{i}^{\prime},\hat{Y}_{i}^{\prime\prime})=\frac{\sum_{snt\in\hat{Y}_{i}^{\prime}}\sum_{b{\text{-}}gram\in snt}\min(C(b{\text{-}}gram,\hat{Y}_{i}^{\prime}|\hat{Y}_{i}^{\prime\prime}),C(b{\text{-}}gram,\hat{Y}_{i}^{\prime\prime}))}{\sum_{\tilde{snt}\in\hat{Y}_{i}^{\prime}}\sum_{b{\text{-}}gram\in\tilde{snt}}C(b{\text{-}}gram,\hat{Y}_{i}^{\prime})}$
@@
 where $snt$ denotes a sentence in an LLM output, $len(\hat{Y})$ denotes the number of tokens in an LLM output, $C(b{\text{-}}gram,\hat{Y}_{i}^{\prime})$ denotes the number of times $b{\text{-}}gram$ appears in $\hat{Y}_{i}^{\prime}$ and $C(b{\text{-}}gram,\hat{Y}_{i}^{\prime}|\hat{Y}_{i}^{\prime\prime})$ denotes the number of times $b{\text{-}}gram$ appears in $\hat{Y}_{i}^{\prime}$ given that it also appears in $\hat{Y}_{i}^{\prime\prime}$ (Papineni et al., 2002; goo, ).
-For symmetry, the minimum of these two BLEU scores for each counterfactual pair is obtained before averaging.
+To achieve symmetry, we use the minimum of these two BLEU scores for each counterfactual pair is obtained before averaging.
 For the same reasons as with CROUGE-L, practitioners should mask protected attribute words in counterfactual output pairs before computing CBLEU.
 Counterfactual Cosine Similarity (CCS).
-Given a sentence transformer $\mathbf{V}:\mathcal{Y}\xrightarrow{}\mathbb{R}^{d}$, this work defines CCS as:
+Given a sentence transformer $\mathbf{V}:\mathcal{Y}\xrightarrow{}\mathbb{R}^{d}$, we define CCS as:
 $CCS=\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{V}(Y_{i}^{\prime})\cdot\mathbf{V}(Y_{i}^{\prime\prime})}{\lVert\mathbf{V}(Y_{i}^{\prime})\rVert\lVert\mathbf{V}(Y_{i}^{\prime\prime})\rVert},$
@@
 To achieve this, these metrics leverage a pre-trained sentiment classifier $Sm:\mathcal{Y}\xrightarrow[]{}[0,1]$.
-This section outlines two counterfactual sentiment metrics: Strict Counterfactual Sentiment Parity, proposed by Huang et al.
-(2020), and an extension of this metric introduced in this work called Weak Counterfactual Sentiment Parity.
+We outline two counterfactual sentiment metrics: Strict Counterfactual Sentiment Parity, proposed by Huang et al.
+(2020), and an extension of this metric called Weak Counterfactual Sentiment Parity.
 [^23] Huang et al.
@@
 Weak Counterfactual Sentiment Parity (WCSP).
-This study presents WCSP, defined as the difference in predicted sentiment rates by a sentiment classifier applied to counterfactually generated LLM output pairs.
+We introduce WCSP, defined as the difference in predicted sentiment rates by a sentiment classifier applied to counterfactually generated LLM output pairs.
 Given a threshold $\tau$ for binarizing sentiment scores, this metric is introduced as follows:
@@
 In practice, practitioners may select an appropriate value of $\tau$ depending on stakeholder values and the sentiment classifier being used.
-3.2 Fairness Metrics for Classification Use Cases
+3.2 Metrics for Classification Use Cases
 It is well-established that classification models can produce unfair outcomes for certain protected attribute groups (Saleiro et al., 2018; Bellamy et al., 2018; Weerts et al., 2023; Feldman et al., 2014; Hardt et al., 2016; Mehrabi et al., 2019).
 Let a classification LLM use case be defined as an LLM tasked with classification, denoted as $\mathcal{M}^{(c)}$, and a population of prompts $\mathcal{P}_{X}$.
-Note that the metrics introduced in this section are confined to binary classification use cases, where $\mathcal{M}^{(c)}:\mathcal{X}\xrightarrow{}\{0,1\},$ with the understanding that evaluating fairness for multiclass classification is a straightforward extension from the binary case (Rouzot et al., 2023).
-For the remainder of this section, assume each prompt in $\mathcal{P}_{X}$ for a given classification LLM use case corresponds to a protected attribute group.
+Here, we present metrics for binary classification use cases, where $\mathcal{M}^{(c)}:\mathcal{X}\xrightarrow{}\{0,1\},$ noting that evaluating fairness for multiclass classification is a straightforward extension from the binary case (Rouzot et al., 2023).
+For the remainder of this section, we assume each prompt in $\mathcal{P}_{X}$ for a given classification LLM use case corresponds to a protected attribute group.
 Under this assumption, traditional machine learning fairness metrics (Bellamy et al., 2018; Saleiro et al., 2018; Weerts et al., 2023) may be applied (Czarnowska et al., 2021).
-Accordingly, these metrics are defined on binary predictions $\hat{Y}_{1},...,\hat{Y}_{N}$, generated from a sample of prompts $X_{1},...,X_{N}\in\mathcal{P}_{X}$, with some metrics also incorporating corresponding ground truth values $Y_{1},...,Y_{N}$.
+Accordingly, we define these metrics on binary predictions $\hat{Y}_{1},...,\hat{Y}_{N}$, generated from a sample of prompts $X_{1},...,X_{N}\in\mathcal{P}_{X}$, with some metrics also incorporating corresponding ground truth values $Y_{1},...,Y_{N}$.
 These metrics effectively assess group fairness (see Section 2.2) between two protected attribute groups $G^{\prime}$ and $G^{\prime\prime}$, with choice of statistical outcome measure $B$ depending on stakeholder values (e.g.
@@
 false positives).
-This section distinguishes between representation fairness metrics, calculated using only predictions, and error-based fairness metrics, calculated using both predictions and ground truth values.
-The formulation of each fairness metric involves calculating the absolute difference between a pair of group-level metrics.
-This calculation yields a range of values between 0 and 1, where smaller values signify a higher level of fairness.
-[^20] Note that while this work considers group fairness in terms of differences between protected attribute groups, as in Gallegos et al.
+We distinguish between representation fairness metrics, calculated using only predictions, and error-based fairness metrics, calculated using both predictions and ground truth values.
+Each fairness metric measures the absolute difference between a pair of group-level metrics.
+This calculation yields a range of values between 0 and 1, where smaller values indicate a higher level of fairness.
+[^24] Note that while we consider group fairness in terms of differences between protected attribute groups, as in Gallegos et al.
 (2023); Bellamy et al.
@@
 3.2.1 Representation Fairness Metrics for Binary Classification
-Representation fairness metrics aim to determine whether protected attribute groups are adequately represented in the positive predictions generated by classifier.
-It is recommended that practitioners reserve this set of metrics for classification LLM use cases for which group-level predicted prevalence rates, i.e.
+Representation fairness metrics aim to determine whether protected attribute groups are adequately represented in the positive predictions generated by a classifier.
+We recommend that practitioners reserve this set of metrics for classification LLM use cases for which group-level predicted prevalence rates, i.e.
 the proportion of predictions belonging to the positive class, should be approximately equal.
 [^25] For example, when predicting which applicants are qualified for a job it may be desirable to target equal rates of positive predictions for males and females.
-However, when predicting certain diseases, this may not be a desirable model behavior.
-A single representation fairness metric, known as Demographic Parity, is defined below.
+However, when predicting certain diseases, for instance, this may not be a desirable model behavior.
+In our framework, we include a single representation fairness metric, Demographic Parity, which measures the absolute difference in group-level predicted prevalence rates.
 Demographic Parity (DP) (Dwork et al., 2011).
@@
 Error-based fairness metrics aim to determine whether disparities in model performance exist across protected attribute groups.
-For error-based fairness, two metrics focused on false negatives, False Negative Rate Difference and False Omission Rate Difference, and two metrics focused on false positives, False Positive Rate Difference and False Discovery Rate Difference, are presented.
+To address error-based fairness, we include two metrics focused on false negatives, False Negative Rate Difference and False Omission Rate Difference, and two metrics focused on false positives, False Positive Rate Difference and False Discovery Rate Difference, in our framework.
 Following Saleiro et al.
-(2018), it is recommended that the choice between these two types of error-based fairness metrics be informed by whether the interventions assigned by the model are assistive (meaning false negatives are undesirable) or punitive (meaning false positives are undesirable) in nature.
+(2018), we recommend that practitioners assess disparities in false negatives (positives) across groups for use cases assigning assistive (punitive) interventions.
+These metrics are defined below.
 False Negative Rate Difference (FNRD) (Bellamy et al., 2018).
@@
 3.2.3 Multiclass Fairness Metrics
-For multiclass classifiers, the framework proposed here follows the fairness guidelines provided by Rouzot et al.
-(2023) and hence recommends conducting class-wise fairness assessments using the appropriate binary classification fairness metrics, as per Sections 3.2.1, 3.2.2, on each of the ‘sensitive’ classes.
+For multiclass classifiers, we follow the fairness guidelines provided by Rouzot et al.
+(2023).
+Hence, we recommend conducting class-wise, one-vs-rest, fairness assessments using the appropriate binary classification fairness metrics, as per Sections 3.2.1, 3.2.2, on each of the ‘sensitive’ classes.
 In particular, Rouzot et al.
 (2023) characterize sensitive classes as outcomes having significant impact on the lives of individuals to whom the model is applied.
-3.3 Fairness Metrics for Recommendation Use Cases
+3.3 Metrics for Recommendation Use Cases
 Zhang et al.
@@
 Specifically, $\mathcal{M}^{(R)}:\mathcal{X}\xrightarrow{}\mathcal{R}^{K}$ maps a prompt $X\in\mathcal{X}$ to an ordered $K$-tuple $\hat{R}\in\mathcal{R}^{K}$ of distinct recommendations from a set of possible recommendations $\mathcal{R}$.
-This section outlines a set of fairness metrics for recommendation LLM use cases, as proposed by Zhang et al.
+We outline a set of fairness metrics for recommendation LLM use cases, as proposed by Zhang et al.
 (2023).
-To ensure consistency with the metrics discussed in Section 3.1.3, this section presents modified versions of these metrics to be pairwise in nature, rather than attribute-wise.
+To maintain consistency with the metrics discussed in Section 3.1.3, we present modified versions of these metrics to be pairwise in nature, rather than attribute-wise.
 Given two protected attribute groups $G^{\prime},G^{\prime\prime},$ and an LLM use case $(\mathcal{M}^{(R)},\mathcal{P}_{X})$, these metrics assess similarity in counterfactually generated recommendation lists.
-Below, each metric is defined according to responses generated from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ which are drawn from $\mathcal{P}_{X|\mathcal{A}}$.
+Below, we define each metric according to responses generated from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ which are drawn from $\mathcal{P}_{X|\mathcal{A}}$.
 In particular, three metrics are presented: Jaccard Similarity, Search Result Page Misinformation Score at K, and Pairwise Ranking Accuracy Gap at K.
@@
 Jaccard Similarity at K (Jaccard-K) (Zhang et al., 2023).
-Below, a pairwise version of Jaccard-K is introduced.
+We present a pairwise version of Jaccard-K.
 This metric calculates the average Jaccard Similarity (Han et al., 2011)—the ratio of the intersection cardinality to the union cardinality—among pairs of counterfactually generated recommendation lists.
@@
 (2021), SERP-K reflects the similarity of two lists, considering both overlap and ranks.
-A modified version of SERP-K, adapted for pairwise application, is introduced as follows:
+We define a modified version of SERP-K, adapted for pairwise application, is introduced as follows:
 $\psi(X_{i}^{\prime},X_{i}^{\prime\prime})=\sum_{v\in\hat{R}_{i}^{\prime}}\frac{I(v\in\hat{R}_{i}^{\prime\prime})*(K-rank(v,\hat{R}_{i}^{\prime})+1)}{K*(K+1)/2},$
@@
 where $\hat{R}_{i}^{\prime},\hat{R}_{i}^{\prime\prime}$ respectively denote the generated lists of recommendations by $\mathcal{M}(X;\theta)$ from the counterfactual input pair $(X_{i}^{\prime},X_{i}^{\prime\prime})$, $v$ is a recommendation from $\hat{R}_{i}^{\prime}$, and $rank(v,\hat{R}_{i}^{\prime})$ denotes the rank of $v$ in $\hat{R}_{i}^{\prime}$.
-Note that the use of $\min(\cdot,\cdot)$ is included to achieve symmetry.
+Note that we use $\min(\cdot,\cdot)$ to achieve symmetry.
 Pairwise Ranking Accuracy Gap at K (PRAG-K) (Zhang et al., 2023).
-Adapted from Beutel et al. (2019), PRAG-K reflects the similarity in pairwise ranking between two recommendation results. A pairwise version of PRAG-K is presented as follows:
+Adapted from Beutel et al. (2019), PRAG-K reflects the similarity in pairwise ranking between two recommendation results. We define a pairwise version of PRAG-K as follows:
 $rankmatch_{i}(v_{1},v_{2})=I(rank(v_{1},\hat{R}_{i}^{\prime})<rank(v_{2},\hat{R}_{i}^{\prime}))*I(rank(v_{1},\hat{R}_{i}^{\prime\prime})<rank(v_{2},\hat{R}_{i}^{\prime\prime}))$
@@
 where $\hat{R}_{i}^{\prime},\hat{R}_{i}^{\prime\prime}$ respectively denote the generated lists of recommendations by $\mathcal{M}(X;\theta)$ from the counterfactual input pair $(X_{i}^{\prime},X_{i}^{\prime\prime})$, $v_{1},v_{2}$ are recommendations from $\hat{R}_{i}^{\prime}$, and $rank(v,\hat{R}_{i})$ denotes the rank of $v$ in $\hat{R}_{i}$.
-As with SERP-K, $\min(\cdot,\cdot)$ is used to achieve symmetry.
-Table 2: Bias and Fairness Evaluation Metrics
+As with SERP-K, we use $\min(\cdot,\cdot)$ to achieve symmetry.
+4 A Unified Framework for Bias and Fairness Assessments of LLM Use Cases
+In general, bias and fairness assessments of LLM use cases do not require satisfying all possible evaluation metrics.
+Instead, practitioners should prioritize and concentrate on a relevant subset of metrics that align with their use case.
+To demystify metric choice for these assessments, we introduce a decision framework that enables practitioners to determine suitable choices of bias and fairness evaluation metrics, drawing inspiration from the classification fairness framework proposed by Saleiro et al.
+(2018).
+Our proposed framework is designed for use cases for which prompts can be sampled from a known population and the task is well-defined.
+We categorize use cases into three distinct groups based on task: 1) text generation and summarization, 2) classification, and 3) recommendation.
+For each category, we map use cases to a set of evaluation metrics to assess the applicable bias and fairness risks.
+This mapping is depicted in Figure 1, and a comprehensive list of bias and fairness evaluation metrics is contained in Table 2.
+First, consider text generation and summarization.
+For this collection of use cases, an important factor in determining relevant bias and fairness metrics is whether the use case upholds FTU, meaning that prompts do not include any mentions of protected attribute words.
+If FTU is not satisfied, we recommend that practitioners include counterfactual fairness and stereotype metrics, as respectively outlined in Sections 3.1.3 and 3.1.2, in their assessments.
+[^26] Note that counterfactual similarity may be too strict in some contexts, particularly in use cases where textual content ought to differ across protected attribute groups (e.g.
+clinical contexts).
+For these use cases, practitioners may opt to focus only on counterfactual sentiment and omit counterfactual similarity metrics from their assessment.
+[^27] Note that stereotype risk, while low, may still exist even if FTU is satisfied.
+Hence, practitioners may wish to conduct stereotype assessments in these scenarios.
+Additionally, we recommend that all text generation and summarization use cases undergo toxicity evaluation, as outlined in Section 3.1.1, regardless of whether or not FTU is upheld.
+For classification use cases, we adopt a modified version of the decision framework proposed by Saleiro et al.
+(2018).
+This framework can be applied to any classification use case where inputs correspond to protected attribute groups.
+Following Saleiro et al.
+(2018), the following approach is recommended: if fairness requires that model predictions exhibit approximately equal predicted prevalence across different groups, representation fairness metrics should be used; Otherwise, error-based fairness metrics should be used.
+For error-based fairness, practitioners should focus on disparities in false negatives (positives), assessed by FNRD and FORD (FPRD and FDRD), if the model is used to assign assistive (punitive) interventions.
+[^28] In the context of fairness, if interventions are punitive, and hence can hurt individuals, it is undesirable for a model to produce false positives disproportionately for any protected attribute group.
+Analogously, having a model produce false negatives disproportionately for any protected attribute group is undesirable in the case of assistive interventions.
+If inputs cannot be mapped to a protected attribute, meaning they are not person-level inputs and they satisfy FTU, then a fairness assessment is not applicable.
+Lastly, for recommendation use cases, counterfactual unfairness is a risk if FTU cannot be satisfied, as shown by Zhang et al.
+(2023).
+Note that counterfactual invariance may not be a desirable property for certain recommendation use cases.
+For instance, it may be preferred to recommend different products for male vs.
+female customers.
+Therefore, if counterfactual invariance is a desired property, we recommend that recommendation use cases not satisfying FTU be assessed for counterfactual unfairness in recommendations using the metrics outlined in Section 3.3.
+Conversely, if a recommendation use case satisfies FTU or if counterfactual invariance is not desired, then a fairness assessment is not applicable.
+Figure 1: Bias and Fairness Evaluation Framework for LLM Use Cases
+5 Experiments
+We conduct a set of experiments using this paper’s companion Python toolkit, LangFair.
+[^29] We use LangFair v0.3.1 for our experiments.
+Given the vast number of studies that have investigated classification fairness (Bellamy et al., 2018; Saleiro et al., 2018; Weerts et al., 2023; Feldman et al., 2014; Zhang et al., 2018; Kamishima et al., 2012) and recommendation fairness (Wang et al., 2023; Li et al., 2023b; Zhang et al., 2023; Beutel et al., 2019), we focus on evaluating bias and fairness for text generation and summarization use cases.
+In particular, we sample from three populations of prompts and use two different LLMs, for a total of six use cases.
+The first two samples, each comprised of 1000 incomplete sentences, are randomly drawn from the RealToxicityPrompts (RTP) dataset (Gehman et al., 2020).
+[^30] In practice, we recommend practitioners adopt the sample sizes used by Wang et al.
+(2024b), ideally sampling 1000 prompts from their use case and generating 25 responses per prompt.
+If obtaining a sample size this large is not feasible, practitioners may generate a larger number of responses per prompt.
+One sample includes prompts with a toxicity level less than 0.2, and the other includes prompts labeled as ‘challenging’.
+Each of these prompts contains an incomplete sentence prepended with instructions to complete the sentence.
+The third sample contains 1000 conversations drawn from the DialogSum (DS) dataset (Chen et al., 2021) prepended with instructions to summarize.
+For the LLMs, we use gemini-1.0-pro and gpt-3.5-turbo-16k.
+To select evaluation metrics, we use our decision framework depicted in Figure 1.
+Since we are dealing with text generation and summarization use cases, we must first determine whether our use cases satisfy FTU.
+Using LangFair’s CounterfactualGenerator class, we parse our three samples of prompts for gender words and find that none of the use cases satisfy FTU.
+We further suppose that counterfactual invariance is required for fairness in our use cases.
+Hence, our recommended assessments are toxicity, stereotype, and counterfactual fairness assessments.
+For all six use cases, we compute the full suite of toxicity, stereotype, and counterfactual metrics presented in Sections 3.1.1, 3.1.2, and 3.1.3, respectively.
+All results are presented in Table 3, and code snippets are contained in Appendix A.
+For each use case, we generate 25 responses for each prompt and compute toxicity and stereotype metrics on the generated responses.
+In the toxicity assessments, we see substantial variation in toxic fraction across prompt sets for the same model and vice versa.
+For instance, toxic fraction is approximately 73 times higher when using gpt-3.5-turbo-16k for sentence completion with the challenging sample of RTP prompts compared to the low-toxicity sample.
+Upon investigation of responses with the highest toxicity scores, we find many highly offensive responses for the RTP use cases.
+Next, we find co-occurrence-based stereotype metrics are far more consistent across use cases, and all values are relatively low in comparison to those found in Bordia and Bowman (2019) and Liang et al.
+(2023).
+However, we find greater variation in stereotype fraction across use cases.
+Manual inspection of responses with highest stereotype scores reveals no cause for concern for the DS use cases, but reveals offensive content for several of the RTP use cases.
+[^31] LangFair provides functionality to compute response-level toxicity, stereotype, and counterfactual scores.
+Lastly, we conduct a counterfactual fairness assessment for gender.
+Among the 1000 prompts sampled from the RTP-challenging, RTP-nontoxic, and DS datasets, 291, 189, and 306 prompts contain gender words, respectively.
+We subset the prompts to retain only prompts mentioning gender words, and create counterfactual input pairs (CIPs) using token-wise substitution.
+For each CIP, we then generate 25 responses per prompt.
+We find far more counterfactual variation for the use cases using gemini-1.0-pro relative to gpt-3.5-turbo-16k.
+Additionally, we find counterfactual responses for DS to be more similar than for either of the RTP prompt samples.
+This is likely due to more opportunity for creativity in sentence completion compared to dialogue summarization.
+After investigating response-level counterfactual scores, we find many instances with large sentiment disparities for the counterfactual responses.
+6 Conclusions
+In this paper, we present an actionable decision framework for selecting bias and fairness evaluation metrics for LLM use cases, introducing several new evaluation metrics as part of the framework.
+This work addresses two gaps in the current literature.
+First, to the best our knowledge, the current literature does not offer a framework for selecting bias and fairness evaluation metrics for LLM use cases.
+Our framework, inspired by Saleiro et al.
+(2018), fills this gap by incorporating use case characteristics and stakeholder values to guide the selection of evaluation metrics.
+Second, our framework tackles limitations of existing LLM bias and fairness evaluation approaches that rely on benchmark datasets containing predefined prompts.
+Instead, our approach uses actual prompts from the practitioner’s use case.
+By considering both prompt-risk and the assigned task of the LLM, our approach provides a more customized risk assessment for the practitioner’s specific use case.
+Furthermore, our proposed framework is highly practical, as all evaluation metrics are computed solely from the LLM output.
+To streamline implementation of the framework, all metrics can be easily computed using this paper’s companion Python toolkit, LangFair.
+Finally, our experiments reveal substantial variation in bias and fairness across use cases, underscoring the importance of conducting these assessments at the use case level.
+7 Limitations
+Despite its strengths, we note two primary limitations of this work.
+First, while our framework aims to encompass the vast majority of use cases for LLMs, we acknowledge that our taxonomy of use cases may not be exhaustive.
+Second, our framework is limited to use cases where prompts are drawn from a known population and does not cater to scenarios where the prompt population is undefined.
+For example, in the context of LLM chatbot applications, it’s unlikely that practitioners can dictate the prompts users input into the chatbot.
+Consequently, our evaluation framework is unable to account for worst-case scenarios in such use cases, where prompts could contain any text input.
+[^32] For more robust bias and fairness monitoring, practitioners may also consider tracking response-level scores.
+Using these scores, practitioners can implement response filtering or targeted human-in-the-loop practices, requiring manual review for any response that yields a concerning score.
+This approach offers advantages over arbitrary human-in-the-loop, where random samples of responses are manually reviewed, and exhaustive human-in-the-loop, which may be infeasible for certain large-scale use cases.
+[^33] Response-level toxicity and stereotype scores can be directly obtained from the respective classifiers.
+For counterfactual fairness, CROUGE-L, CBLEU, and CCS may be computed at the response-pair level and disparities in sentiment scores can be calculated using differences or ratios of sentiment-classifier outputs.
+These approaches can be implemented with the LangFair library.
+Acknowledgements
+We wish to thank Mohit Singh Chauhan, Blake Aber, Piero Ferrante, Xue (Crystal) Gu, Almira Pillay, Zeya Ahmad, Kee Siong Ng, Huiwen Hu, and Vasistha Singhal Vinod for their helpful suggestions as well as David Skarbrevik and Viren Bajaj for their contributions to the LangFair library.
+Appendix A Code Snippets
+Below, we provide code snippets from our experiments using LangFair. Note that the printed metric values are purely illustrative.
+# Construct LangChain LLM
+from langchain_google_vertexai import ChatVertexAI
+from langchain_core.rate_limiters import InMemoryRateLimiter
+rate_limiter = InMemoryRateLimiter(
+requests_per_second=4.5, check_every_n_seconds=0.5, max_bucket_size=280,
+llm = ChatVertexAI(
+model_name="gemini-pro", temperature=0.3, rate_limiter=rate_limiter
+# Generate 25 LLM responses per prompt
+from langfair.generator import ResponseGenerator
+rg = ResponseGenerator(langchain_llm=llm)
+generations = await rg.generate_responses(prompts=prompts, count=25)
+responses = generations["data"]["response"]
+duplicated_prompts = generations["data"]["prompt"] # so prompts correspond to responses
+# Compute toxicity metrics
+import torch
+from langfair.metrics.toxicity import ToxicityMetrics
+device = torch.device("cuda")
+tm = ToxicityMetrics(device=device)
+tox_result = tm.evaluate(
+prompts=duplicated_prompts,
+responses=responses,
+return_data=True
+tox_result["metrics"]
+# # Output is below
+# {’Toxic Fraction’: 0.0004,
+# ’Expected Maximum Toxicity’: 0.013845130120171235,
+# ’Toxicity Probability’: 0.01}
+# Compute stereotype metrics
+from langfair.metrics.stereotype import StereotypeMetrics
+sm = StereotypeMetrics()
+stereo_result = sm.evaluate(responses=responses, categories=["gender"])
+stereo_result["metrics"]
+# # Output is below
+# {’Stereotype Association’: 0.3172750176745329,
+# ’Cooccurrence Bias’: 0.44766333654278373,
+# ’Stereotype Fraction - gender’: 0.08}
+# Check for FTU
+from langfair.generator.counterfactual import CounterfactualGenerator
+cg = CounterfactualGenerator(langchain_llm=llm)
+ftu_result = cg.check_ftu(
+prompts=prompts,
+attribute="gender",
+subset_prompts=True
+pd.DataFrame(ftu_result["data"])
+# Generate counterfactual responses
+cf_generations = await cg.generate_responses(
+prompts=prompts, attribute="gender", count=25
+male_responses = cf_generations["data"]["male_response"]
+female_responses = cf_generations["data"]["female_response"]
+# Compute counterfactual metrics
+from langfair.metrics.counterfactual import CounterfactualMetrics
+cm = CounterfactualMetrics()
+cf_result = cm.evaluate(
+texts1=male_responses,
+texts2=female_responses,
+attribute="gender"
+cf_result["metrics"]
+# # Output is below
+# {’Cosine Similarity’: 0.8318708,
+# ’RougeL Similarity’: 0.5195852482361165,
+# ’Bleu Similarity’: 0.3278433712872481,
+# ’Sentiment Bias’: 0.0009947145187601957}
+Appendix B Supplemental Tables
+Table 1: Taxonomy of LLM Use Cases and Associated Bias/Fairness Risks
+Use Case Category Description Examples Bias/Fairness Risk
+Text Generation and Summarization LLM generates text outputs that are not constrained to a predefined set of classes or list elements Create personalized outreach messages to individuals; Summarize clinical notes Toxic text, stereotypes*, counterfactual fairness*
+Classification LLM classifies a text input among a pre-defined set of classes Classify intent of customer support inquiries to assign assistance; Classify customer feedback as positive or negative to assign follow-ups Allocational harms**
+Recommendation LLM generates lists of recommendations Generate lists of recommended products; Generate lists of recommended news articles Counterfactual fairness*
+*Risk is applicable if FTU is not satisfied. Counterfactual fairness may not be relevant in certain contexts.
+**Risk is applicable if text inputs correspond to a protected attribute.
+Table 2: Glossary of Bias and Fairness Evaluation Metrics
 Evaluation Metric Required Input
@@
    PRAG-K Counterfactual pairs of generated recommendation lists of length $K$
-4 A Unified Framework for Bias and Fairness Assessments of LLM Use Cases
-In general, bias and fairness assessments of LLM use cases do not require satisfying all possible evaluation metrics.
-Instead, practitioners should prioritize and concentrate on a relevant subset of metrics that align with their use case.
-To demystify metric choice for these assessments, this section introduces a decision framework that enables practitioners to determine suitable choices of bias and fairness evaluation metrics, drawing inspiration from the classification fairness framework proposed by Saleiro et al.
-(2018).
-The scope of this framework is restricted to use cases for which prompts can be sampled from a known population and the task is well-defined.
-Use cases are categorized into three distinct groups based on task: 1) text generation and summarization, 2) classification, and 3) recommendation.
-The framework includes suitable metrics for each category to assess the potential bias and fairness risks that align with the specific characteristics of the use case.
-The mapping of these metrics is illustrated in Figure 1, and a thorough discussion of this mapping is provided below.
-For a comprehensive list of bias and fairness evaluation metrics, the reader may refer to Table 2.
-First, consider text generation and summarization.
-For this collection of use cases, an important factor in determining relevant bias and fairness metrics is whether the use case upholds FTU, meaning that prompts do not include any mentions of protected attribute words.
-If FTU cannot be achieved for this type of use case, it is recommended that practitioners include counterfactual fairness and stereotype metrics, as respectively outlined in Sections 3.1.3 and 3.1.2, in their assessments.
-[^22] Note that counterfactual similarity may be too strict in some contexts, particularly in scenarios where textual content ought to differ across protected attribute groups (e.g.
-clinical contexts).
-For these scenarios, practitioners may opt to focus only on counterfactual sentiment and omit counterfactual similarity metrics from their assessment.
-[^23] Note that stereotype risk, while low, may still exist even if FTU is satisfied.
-Hence, practitioners may wish to conduct stereotype assessments in these scenarios.
-Additionally, it is recommended that all text generation and summarization use cases undergo toxicity evaluation, as outlined in Section 3.1.1, regardless of whether or not FTU is achieved.
-For classification use cases, a modified version of the decision framework proposed by Saleiro et al.
-(2018) is adopted.
-This framework can be applied to any classification use case where inputs correspond to protected attribute groups.
-Following Saleiro et al.
-(2018), the following approach is recommended: if fairness necessitates that model predictions exhibit approximately equal predicted prevalence across different groups, representation fairness metrics should be used; Otherwise, error-based fairness metrics should be used.
-For error-based fairness, practitioners should focus on disparities in false negatives (positives), assessed by FNRD and FORD (FPRD and FDRD), if the model is used to assign assistive (punitive) interventions.
-[^24] In the context of fairness, if interventions are punitive, and hence can hurt individuals, it is undesirable for a model to produce false positives disproportionately for any protected attribute group.
-Analogously, having a model produce false negatives disproportionately for any protected attribute group is undesirable in the case of assistive interventions.
-If inputs cannot be mapped to a protected attribute, meaning they are not person-level inputs and they satisfy FTU, then a fairness assessment is not applicable.
-Third, for recommendation use cases, counterfactual unfairness is a risk if FTU cannot be satisfied, as shown by Zhang et al.
-(2023).
-Note that counterfactual invariance may not be a desirable property for certain recommendation use cases.
-For instance, it may be preferred to recommend different products for male vs.
-female customers.
-Hence, if counterfactual invariance is a desired property, it is recommended that recommendation use cases not satisfying FTU be assessed for counterfactual unfairness in recommendations using the metrics outlined in Section 3.3.
-On the other hand, if a recommendation use case satisfies FTU or if counterfactual invariance is not desired, then a fairness assessment is not applicable.
-It is important to note that the framework presented here is intended for use cases in which the volume of LLM responses makes exhaustive human review infeasible.
-Accordingly, it is important to note that in situations where the practitioner manually reviews each generated output, the proposed evaluations may be unnecessary if the concerns regarding bias and fairness can be effectively addressed by the output reviewer themselves.
-Figure 1: Bias and Fairness Evaluation Framework for LLM Use Cases
-5 Conclusion
-This paper proposes an actionable decision framework for selecting bias and fairness evaluation metrics for LLM use cases, introducing several new evaluation metrics as part of the framework.
-This work addresses two gaps in the current literature.
-First, to the best of the author’s knowledge, the current literature does not offer a framework for selecting bias and fairness evaluation metrics for LLM use cases.
-To fill this gap, the proposed framework draws inspiration from Saleiro et al.
-(2018) and incorporates use case characteristics and stakeholder values to guide the selection of evaluation metrics.
-Second, this framework tackles limitations of existing LLM bias and fairness evaluation approaches that rely on benchmark data sets containing predefined prompts.
-Instead, the approach outlined in this work involves using actual prompts from the practitioner’s use case.
-By considering both prompt-risk and the assigned task of the LLM, this approach provides a more customized risk assessment for the practitioner’s specific use case.
-Furthermore, the proposed framework aims to enhance practicality and ease of implementation, as all evaluation metrics are computed solely from the LLM output.
-Acknowledgements
-Thank you to Mohit Singh Chauhan, Blake Aber, Piero Ferrante, Xue (Crystal) Gu, Almira Pillay, Zeya Ahmad, and Vasistha Singhal Vinod for your helpful suggestions.
+Table 3: Bias and Fairness Evaluation Results: Text Generation / Summarization Experiments
+GPT-3.5-Turbo Gemini-1.0-Pro
+Metric RTP-Challenging RTP-Nontoxic Dialogue-Sum RTP-Challenging RTP-Nontoxic Dialogue
+Toxicity Metrics
+   Toxic Fraction 0.437 0.006 0.003 0.158 0.004 0.000
+   Expected Maximum Toxicity 0.547 0.021 0.006 0.588 0.050 0.005
+   Toxicity Probability 0.578 0.018 0.005 0.734 0.048 0.002
+   Number of responses 25000 25000 25000 25000 25000 25000
+Stereotype Metrics
+   Stereotype Association 0.394 0.402 0.334 0.337 0.302 0.317
+   Cooccurrence Bias 0.647 0.867 0.533 0.845 0.739 0.537
+   Stereotype Fraction - gender 0.148 0.073 0.213 0.125 0.033 0.140
+   Number of responses 25000 25000 25000 25000 25000 25000
+Counterfactual Metrics
+   Cosine Similarity 0.705 0.692 0.912 0.545 0.521 0.801
+   ROUGE-L Similarity 0.616 0.587 0.656 0.299 0.309 0.467
+   BLEU Similarity 0.465 0.421 0.459 0.192 0.163 0.270
+   Strict Sentiment Bias 0.007 0.003 0.001 0.014 0.002 0.003
+   Number of response pairs* 4501 4533 7611 7049 3850 7650
+*Counterfactual metrics were computed using 25 counterfactual LLM responses per counterfactual input pair.
+We constructed 291, 189, and 306 counterfactual input pairs for the RTP-Challenging, RTP-Nontoxic, and Dialogue-Sum datasets, respectively.
+Note that response pairs were excluded if either of the responses were blocked by content filters.
+As a result, the total number of response pairs used was less than 25 times the number of counterfactual input pairs.